### v1

# Xing Huanga), Panpan Renb), Feng-Yu Wanga)

Abstract
Let $\tilde{W}^{-\delta,k}$ be the local negative Sobolev space on $\mathbb{R}^{d}$ with indexes $\delta\in[0,\infty)$ and $k\in[1,\infty].$ We study McKean-Vlasov SDEs with interaction kernels in $\tilde{W}^{-\delta,k}.$ By developing a time-shift argument which allows the singular interactions vanishing at time 0, the global well-posedness is proved for regular enough initial distributions and any singular indexes $(\delta,k)\in[0,\infty)\times[1,\infty],$ and for any initial distributions provided $\delta+\frac{d}{k}<1$. Moreover, the relative entropy and the $\|\cdot\|_{\delta,k*}$-distance induced by $\tilde{W}^{-\delta,k}$ are estimated for the time-marginal distributions of solutions by using the Wasserstein distance of initial distributions, which describe the regularity of the solution in initial distribution. In particular, the main results apply to Nemytskii-type SDEs which point-wisely depend on the density function and its derivatives, as well as McKean-Vlasov SDEs with interactions more singular than Riesz kernels.
AMS Subject Classification: 60H10, 60H50.
Keywords: McKean-Vlasov SDE, local distributional interaction, $\|\cdot\|_{\delta,k*}$-distance,
entropy-cost inequality.
1 Introduction
Let $\mathscr{P}$ be the set of all probability measures on $\mathbb{R}^{d}$ equipped with the weak topology. Consider the following McKean-Vlasov SDE on $\mathbb{R}^{d}$:
(1.1) $\text{\rm{d}}X_{t}=b_{t}(X_{t},\mathscr{L}_{X_{t}})\text{\rm{d}}t+\text{\rm{d}}W_{t},\ \ t\geq 0,$
where $(W_{t})_{t\geq 0}$ is a $d$-dimensional Brownian motion on a probability base (i.e. complete filtered probability space) $(\Omega,\{\mathscr{F}_{t}\}_{t\geq 0},\mathscr{F},\mathbb{P})$, $\mathscr{L}_{X_{t}}$ is the distribution of $X_{t}$, $\tilde{\mathscr{P}}$ is a measurable subspace of $\mathscr{P}$ to be determined by the singularity of $b_{t}(x,\mu)$ in $\mu$, and
$b:[0,\infty)\times\mathbb{R}^{d}\times\tilde{\mathscr{P}}\rightarrow\mathbb{R}^{d}$
is measurable.
We are interested in the case with singular interaction kernels where the drift includes
(1.2) $b_{t}(x,\mu):=\int_{\mathbb{R}^{d}}h_{t}(x-y)\mu(\text{\rm{d}}y),$
for $h_{t}$ belonging to a local negative Sobolev space, and the integral with respect to $\mu$ is understood as duality. More precisely, for any $\delta\in[0,\infty)$ and $k\in[1,\infty],$ let $\tilde{W}^{-\delta,k}(\mathbb{R}^{d};\mathbb{R}^{d})$ be the closure of $\mathscr{B}_{b}(\mathbb{R}^{d};\mathbb{R}^{d})$ under the norm
$\|f\|_{\tilde{W}^{-\delta,k}}:=\sup_{z\in\mathbb{R}^{d}}\big\|1_{B(z,1)}(1-\Delta)^{-\frac{\delta}{2}}f\big\|_{L^{k}}$
for $B(z,1):=\{x\in\mathbb{R}^{d}:|x-z|\leq 1\}$ and $\|\cdot\|_{L^{k}}$ being the $L^{k}$-norm with respect to the Lebesgue measure. Let $\tilde{W}_{*}^{-\delta,k}$ be the dual space of $\tilde{W}^{-\delta,k}.$ We study (1.1) for $h_{t}\in\tilde{W}^{-\delta,k}$ for some $(\delta,k)\in[0,\infty)\times[1,\infty],$ where $b_{t}$ in (1.2) is well-defined as duality for
$\mu\in\mathscr{P}_{\delta,k*}:=\mathscr{P}\cap\tilde{W}_{*}^{-\delta,k}:=\left\{\mu\in\mathscr{P}:\ \|\mu\|_{\delta,k*}:=\sup_{f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\|f\|_{\tilde{W}^{-\delta,k}}\leq 1}|\mu(f)|<\infty\right\}.$
It is easy to see that $\mathscr{P}_{\delta,k*}$ is a closed subset of $\tilde{W}^{-\delta,k}_{*},$ where $\tilde{W}^{-\delta,k}_{*}$ is a Banach space with norm
$\|\mu-\nu\|_{\delta,k*}:=\sup_{f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\|f\|_{\tilde{W}^{-\delta,k}}\leq 1}\big|\mu(f)-\nu(f)\big|,\ \ \ \mu,\nu\in\tilde{W}^{-\delta,k}_{*}.$
When $\delta=0$, we denote
$\big(\tilde{W}^{-0,k}(\mathbb{R}^{d};\mathbb{R}^{d}),\|\cdot\|_{\tilde{W}^{-0,k}}\big)=\big(\tilde{L}^{k}(\mathbb{R}^{d};\mathbb{R}^{d}),\|\cdot\|_{\tilde{L}^{k}}\big).$
Similarly, let $\tilde{W}^{-\delta,k}$ be the closure of $\mathscr{B}_{b}(\mathbb{R}^{d})$ under $\|\cdot\|_{\tilde{W}^{-\delta,k}}$, where $\mathscr{B}_{b}(\mathbb{R}^{d})$ is the space of all bounded measurable functions on $\mathbb{R}^{d}$, and denote
$\big(\tilde{W}^{-0,k},\|\cdot\|_{\tilde{W}^{-0,k}}\big)=\big(\tilde{L}^{k},\|\cdot\|_{\tilde{L}^{k}}\big).$
Typical examples for $h_{t}\in\tilde{L}^{k}(\mathbb{R}^{d};\mathbb{R}^{d})$ include Riesz kernels satisfying
$|h_{t}(z)|\leq\frac{c}{|z|^{\beta}},\ \ 0\neq z\in\mathbb{R}^{d}$
for some constants $c>0$ and $\beta\in(0,d)$, which belongs to $\tilde{L}^{k}(\mathbb{R}^{d};\mathbb{R}^{d})$ for $k\in[1,\frac{d}{\beta})$. In this case, the well-posedness and regularity estimates have been established in our recent paper [8] by establishing entropy-cost inequalities, which describe the regularity of time-marginal distributions of solutions with respect to initial distributions. However, these regularity estimates remain open when $h_{t}$ is merely distributional, although the well-posedness has been intensively studied in [1, 4, 9] when $h_{t}$ belongs to a suitable Besov space, where the noise can be $\alpha$-stable and/or degenerate. See also [2, 3, 4, 5] for the weak well-posedness of SDEs with distributional drifts without interaction (i.e. distribution independent), for which the entropy-cost inequality remains open.
In this paper, we prove the well-posedness and establish regularity estimates for local distributional interaction kernels with arbitrary singular indexes $(\delta,k)\in[0,\infty)\times[1,\infty]$, where the well-posedness result is also new in the literature, see Remark 2.2 below.
To solve (1.1) for $b$ in (1.2) with $h_{t}\in\tilde{W}^{-\delta,k}(\mathbb{R}^{d};\mathbb{R}^{d}),$ we first present some a-priori estimates on the time-marginal distribution of the solution, in particular we need to verify that $\mathscr{L}_{X_{t}}\in\mathscr{P}_{\delta,k*},$ which ensures that the drift in (1.2) is well-defined for $\mu=\mathscr{L}_{X_{t}}.$ To this end, we introduce the following estimate for the standard heat semigroup $P_{t}^{0}$ of the Brownian motion $W_{t}$ (see Lemma 3.1 below): there exists an increasing function $B:[0,\infty)\rightarrow(0,\infty)$ such that
(1.3) $\begin{split}&\|\nabla^{i}P_{t}^{0}\|_{\tilde{W}^{-\delta,k}\rightarrow\tilde{W}^{-\varepsilon,p}}:=\sup_{\|f\|_{\tilde{W}^{-\delta,k}}\leq 1}\|P_{t}^{0}f\|_{\tilde{W}^{-\varepsilon,p}}\leq B_{\delta-\varepsilon}t^{-\frac{i+\delta-\varepsilon}{2}-\frac{d(p-k)}{2pk}},\\ &\qquad\ \ t>0,\ i=0,1,\ \ \infty>\delta\geq\varepsilon\geq 0,\ \ \infty\geq p\geq k\geq 1.\end{split}$
When $i=0,$ this suggests that the time-marginal distribution $(\mu_{t}:=\mathscr{L}_{X_{t}})_{t\in[0,T]}$ of solution to (1.1) satisfies
$\rho_{\varepsilon,p;\delta,k}^{T}(\mu):=\sup_{t\in[0,T]}t^{\frac{\delta-\varepsilon}{2}+\frac{d(p-k)}{2pk}}\|\mu_{t}\|_{\delta,k*}\leq C(T)\|\mu_{0}\|_{\varepsilon,p*},\ \ \ T\in(0,\infty)$
for some increasing $C:(0,\infty)\rightarrow(0,\infty)$. Thus, given initial value $X_{0}$ with $\mathscr{L}_{X_{0}}\in\mathscr{P}_{\varepsilon,p*},$ it is reasonable to solve (1.1) up to time $T$ with time-marginal distributions belonging to the path space
(1.4) $\mathscr{C}_{\varepsilon,p;\delta,k}^{T}:=\Big\{\mu\in C^{w}([0,T];\mathscr{P}):\ \rho^{T}_{\varepsilon,p;\delta,k}(\mu)<\infty\Big\},$
where $C^{w}([0,T];\mathscr{P})$ is the set of all weakly continuous maps from $[0,T]$ to $\mathscr{P}.$ This observation leads to the following notion of $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution of (1.1).
Definition 1.1 (Maximal $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution).
Let $1\leq k\leq p\leq\infty$, $0\leq\varepsilon\leq\delta<\infty$, and $\tilde{\mathscr{P}}=\mathscr{P}_{\delta,k*}$.
  1. (1)
    We call $(X_{t})_{t\in[0,\tau)}$ a maximal strong $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution of (1.1) with life time $\tau$, if $\mathscr{L}_{X_{0}}\in\mathscr{P}_{\varepsilon,p*}$, $\tau\in(0,\infty]$ such that
    $\limsup_{t\uparrow\tau}\|\mathscr{L}_{X_{t}}\|_{\delta,k*}=\infty\ \ \text{when }\ \tau<\infty,$
    $(\mathscr{L}_{X_{t}})_{t\in[0,T]}\in\mathscr{C}_{\varepsilon,p;\delta,k}^{T}$ for any $T\in(0,\tau)$, and $\mathbb{P}$-a.s.
    $X_{t}=X_{0}+\int_{0}^{t}b_{s}(X_{s},\mathscr{L}_{X_{s}})\text{\rm{d}}s+W_{t},\ \ t\in[0,\tau).$
    When $\tau=\infty$, we call $(X_{t})_{t\geq 0}$ a global strong $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution of (1.1). For any $T\in(0,\tau)$, we call $(X_{t},W_{t})_{t\in[0,T]}$ a strong $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution of (1.1) up to time $T$.
  2. (2)
    A couple $(X_{t},W_{t})_{t\in[0,\tau)}$ is called a maximal weak $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution of (1.1) with initial distribution $\gamma\in\mathscr{P}_{\varepsilon,p*}$, if there exists a probability base $(\Omega,\{\mathscr{F}_{t}\}_{t\in[0,\tau)},\mathscr{F},\mathbb{P})$ such that $(W_{t})_{t\in[0,\tau)}$ is a $d$-dimensional Brownian motion, $\mathscr{L}_{X_{0}}=\gamma$ and $(X_{t})_{t\in[0,\tau)}$ is a maximal strong $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution of (1.1). For any $T\in(0,\tau)$, $(X_{t},W_{t})_{t\in[0,T]}$ is called a weak $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution of (1.1) up to time $T$.
  3. (3)
    If (1.1) has a maximal weak $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution with initial distribution $\gamma$, and any two maximal weak $\mathscr{C}_{\varepsilon,p;\delta,k}$-solutions with initial distribution $\gamma$ have common life time and distribution, then we say that (1.1) has a unique maximal weak $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution with initial distribution $\gamma$. In this case, we denote the life time by $\tau(\gamma)$, and set
    $P_{t}^{*}\gamma:=\mathscr{L}_{X_{t}},\ \ t\in[0,\tau(\gamma)).$
In Section 2, we state the main results of the paper on the well-posedness and regularity estimates for $\mathscr{C}_{\varepsilon,p;\delta,k}$-solutions of (1.1), see Theorem 2.1 and Theorem 2.3. Section 3 contains necessary preparations, which will be used in Sections 4 and 5 to prove these two theorems respectively.
2 Main results
To solve (1.1), we make the following assumptions where the drift is Lipschitz continuous in distribution under the $\|\cdot\|_{\delta,k*}$ distance. To cancel the singularity in small times caused by (1.3), we allow the drift decaying with rate $t^{\tau}$ for some $\tau>0$, which can be dropped (i.e. $\tau=0$) by a time-shift argument of the drift as shown in the proof of Theorem 2.1(3).
  1. (A)
    Let $1\leq k\leq\infty$ and $\delta,\kappa\in[0,\infty)$. There exists increasing $K:(0,\infty)\rightarrow[1,\infty)$ such that
    $|b_{t}(x,\nu)|\leq K_{t}t^{\kappa}\|\nu\|_{\delta,k*},\ \ \ \ |b_{t}(x,\mu)-b_{t}(x,\nu)|\leq K_{t}t^{\kappa}\|\mu-\nu\|_{\delta,k*}$
    hold for any $t\in(0,\infty),\ x\in\mathbb{R}^{d},\ \mu,\nu\in\mathscr{P}_{\delta,k*}.$
For any $r>0$ and $\mu\in\mathscr{P}$, let $\mu*N_{0,r}\in\mathscr{P}$ be the distribution of the Brownian motion at time $r$ with initial distribution $\mu$, i.e.
$(\mu*N_{0,r})(A):=(2\pi r)^{-\frac{d}{2}}\int_{\mathbb{R}^{d}\times A}\text{\rm{e}}^{-\frac{|x-y|^{2}}{2r}}\mu(\text{\rm{d}}x)\text{\rm{d}}y,\ \ A\in\mathscr{B}(\mathbb{R}^{d}).$
Let
$\hat{\mathscr{P}_{r}}:=\big\{\mu*N_{0,r}:\ \mu\in\mathscr{P}\big\},\ \ \ \ r>0,$
$\hat{\mathscr{P}_{0}}:=\cup_{r>0}\big\{\mu*N_{0,r}:\ \mu\in\mathscr{P}\big\}.$
Under (A), let $\varepsilon\in[0,\delta]$ and $p\in[k,\infty]$ such that
(2.1) $\eta:=\delta-\varepsilon+\frac{d(p-k)}{pk}<1+2\kappa.$
In this case,
(2.2) $\theta:=\frac{2}{1-(\eta-2\kappa)^{+}}\in[2,\infty).$
The following result ensures the global well-posedness of (1.1) for initial distributions in $\hat{\mathscr{P}_{0}}$, and for arbitrary initial distributions when $\varepsilon=0$ and $p=\infty$. Moreover, when $\varepsilon>0$ or $p<\infty$, (1.1) has a unique maximal $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution for initial distributions in $\mathscr{P}_{\varepsilon,p*}\setminus\hat{\mathscr{P}_{0}}$.
Theorem 2.1.
Assume (A) and let $\varepsilon\in[0,\delta]$ and $p\in[k,\infty]$ satisfying (2.1).
  1. $(1)$
    For any $\mathscr{F}_{0}$-measurable initial value $X_{0}$ with $\gamma:=\mathscr{L}_{X_{0}}\in\mathscr{P}_{\varepsilon,p*}$, (1.1) has a unique maximal (weak and strong) $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution, and there exists increasing
    $C_{\gamma}:[1,\infty)\times(0,\tau(\gamma))\rightarrow(0,\infty)$
    such that
    (2.3) $\mathbb{E}\bigg[\sup_{s\in[0,t]}|X_{s}|^{q}\bigg|\mathscr{F}_{0}\bigg]\leq C_{\gamma}(q,t)(1+|X_{0}|^{q}),\ \ q\in[1,\infty),\ t\in(0,\tau(\gamma)).$
    If $\varepsilon=0$ and $p=\infty$, then $\tau(\gamma)=\infty$ and $C_{\gamma}(q,t)=C(q,t)$ is independent of $\gamma\in\mathscr{P}.$
  2. $(2)$
    Let $\theta$ be in (2.2). For any $n\in\mathbb{N}$, there exist constants $A_{n}\in(0,\infty)$ and $\lambda_{n}\in[0,\infty)$ such that for any $\gamma\in\mathscr{P}_{\varepsilon,p*}$,
    (2.4) $\tau(\gamma)>\tau_{n}(\gamma):=\begin{cases}n,&\text{if}\ \varepsilon=0,p=\infty,\\ \min\big\{n,\ \big(A_{n}\|\gamma\|_{\varepsilon,p*}^{\theta}\text{\rm{e}}^{A_{n}\|\gamma\|_{\varepsilon,p*}^{\theta}}\big)^{-1}\big\},&\text{otherwise},\end{cases}$
    (2.5) $\sup_{t\in(0,\tau_{n}(\gamma)]}t^{\frac{\eta}{2}}\text{\rm{e}}^{-\lambda_{n}t}\|P_{t}^{*}\gamma\|_{\delta,k*}\leq 2B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*},$
    where $\lambda_{n}=0$ if $\tau_{n}(\gamma)<n$. In particular, if $\varepsilon=0$ and $p=\infty$, then $\tau(\gamma)=\infty$ for any $\gamma\in\mathscr{P}$, and
    $\sup_{t\in(0,T],\gamma\in\mathscr{P}}t^{\frac{\eta}{2}}\|P_{t}^{*}\gamma\|_{\delta,k*}<\infty,\ \ T\in(0,\infty).$
  3. $(3)$
    For any $\mathscr{F}_{0}$-measurable initial value $X_{0}$ with $\gamma:=\mathscr{L}_{X_{0}}\in\hat{\mathscr{P}_{0}}$, the SDE (1.1) has a unique global (weak and strong) $\mathscr{C}_{\delta,k;\delta,k}$-solution with life time $\tau(\gamma)=\infty$, and
    (2.6) $\sup_{\gamma\in\hat{\mathscr{P}_{r}},s\in[0,t]}\|P_{s}^{*}\gamma\|_{\delta,k*}\leq c_{r}(t),\ \ \ r,t\in(0,\infty)$
    holds for some $c_{r}(t)\in(0,\infty)$ increasing in $t$ but decreasing in $r$.
Remark 2.2.
Let $b$ be in (1.2) such that
$\|h_{t}\|_{\tilde{W}^{-\delta,k}}\leq K_{t}t^{\kappa},\ \ \ t\geq 0.$
Then assertions in Theorem 2.1 hold for $\varepsilon\in[0,\delta]$ and $p\in[k,\infty]$ satisfying (2.1).
When $\kappa=0$, the condition (2.1) coincides with [4, (1.33)] for $\alpha=2,q_{b}=\infty,p_{0}=\frac{p}{p-1},\rho_{0}=k,\beta_{0}=\varepsilon$ and $\beta_{b}=-\delta$. In this case, [4, Theorem 1.9] ensures the well-posedness and density estimates for (1.1) provided $\|h_{t}\|_{\mathbb{B}_{k,\infty}^{\delta}}<\infty$. If moreover $p=\infty$ and $\varepsilon=0$, (2.1) becomes $\delta<1-\frac{d}{k},$ so that [1, Theorem 1] with $\alpha=2$ implies the global well-posedness of (1.1) for $\|h\|_{\mathbb{B}_{k,\infty}^{-\delta}}<\infty.$ These results do not cover Theorem 2.1 since $\|h\|_{\mathbb{B}_{k,\infty}^{-\delta}}<\infty$ may fail for $h\in\tilde{W}^{-\delta,k}(\mathbb{R}^{d};\mathbb{R}^{d}).$
For $\delta\in[0,\infty)$ and $k\in[1,\infty]$, let $W^{-\delta,k}(\mathbb{R}^{d};\mathbb{R}^{d})$ be the closure of $C_{0}^{\infty}(\mathbb{R}^{d};\mathbb{R}^{d})$ with respect to the negative Sobolev norm $\|f\|_{W^{-\delta,k}}:=\|(1-\Delta)^{-\frac{\delta}{2}}f\|_{L^{k}}$. The propagation of chaos is derived in [10] for $h_{t}\in W^{-1,\infty}(\mathbb{R}^{d};\mathbb{R}^{d}).$ Since $\tilde{W}^{-\delta,k}(\mathbb{R}^{d};\mathbb{R}^{d})$ is larger than $W^{-\delta,k}(\mathbb{R}^{d};\mathbb{R}^{d})$, the propagation of chaos remains open for local distributional interactions in $W^{-\delta,k}(\mathbb{R}^{d};\mathbb{R}^{d})$ for $\delta>0$ and $k\geq 1$.
Moreover, Theorem 2.1(3) provides the global well-posedness for the $\mathscr{C}_{\delta,k;\delta,k}$-solution of (1.1) with arbitrary singular indexes $\delta\in[0,\infty)$ and $k\in[1,\infty]$. This is new in the literature, which in particular applies to interactions more singular than Riesz/Coulomb kernels [11, 14], and also applies to density-derivative dependent (i.e. Nemytskii-type) SDEs, see Example $2.4$ and Example $2.5$ below.
2.1 Regularity estimates
[…]
${\rm Ent}(\gamma|\tilde{\gamma}):=\begin{cases}\gamma\big(\log\frac{\text{\rm{d}}\gamma}{\text{\rm{d}}\tilde{\gamma}}\big),\ &\text{if}\ \frac{\text{\rm{d}}\gamma}{\text{\rm{d}}\tilde{\gamma}}\ \text{exists},\\ \infty,\ &\text{otherwise},\end{cases}$
and for any constant $q\in[1,\infty)$,
$\mathbb{W}_{q}(\gamma,\tilde{\gamma}):=\inf_{\pi\in\mathscr{C}(\gamma,\tilde{\gamma})}\bigg(\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{q}\pi(\text{\rm{d}}x,\text{\rm{d}}y)\bigg)^{\frac{1}{q}},$
[…]
(2.7) $k_{t}(\gamma):=\|\gamma\|_{\varepsilon,p*}\lor\Big(\sup_{s\in(0,t]}s^{\frac{\delta-\varepsilon}{2}+\frac{d(p-k)}{2pk}}\|P_{s}^{*}\gamma\|_{\delta,k*}\Big),\ \ \ t\in(0,\tau(\gamma)),\ \gamma\in\mathscr{P}_{\varepsilon,p*}.$
Let $\theta$ be in (2.2). For any $\gamma,\tilde{\gamma}\in\mathscr{P}_{\varepsilon,p*}$ and increasing function $\beta:(0,\infty)\rightarrow(0,\infty)$, let
$),K_{t,\beta}^{(\theta)}(\gamma,\tilde{\gamma}):=\exp\Big[\beta_{t}\text{\rm{e}}^{\beta_{t}(tk_{t}(\gamma)^{\theta}+tk_{t}(\tilde{\gamma})^{\theta})}\Big],\ \ \ t\in(0,\tau(\gamma)\land\tau(\tilde{\gamma}$)),
(2.8) $s_{t}(\theta^{\prime},\gamma):=t\land[k_{t}(\gamma)^{-\theta^{\prime}}],\ \ \ \theta^{\prime}\in(\theta,\infty),\ t\in\big(0,\tau(\gamma)\big).$
Theorem 2.3.
Assume (A), let $p\in[k,\infty]$ and $\varepsilon\in[0,\delta\land\frac{d(p-1)}{p}]$ such that
(2.9) $\eta:=\delta-\varepsilon+\frac{d(p-k)}{pk}<1\lor(\frac{1}{2}+\kappa),\ \ \ \delta<1\land\Big(2-\frac{d}{k}\Big)+(2\kappa-\eta)^{+}.$
Then $1+(2\kappa-\eta)^{+}-\eta>(\varepsilon+\frac{d}{p}-\frac{d}{k})^{+}$, and for any $q\in[1,\infty)$ satisfying
(2.10) $\frac{\varepsilon+\frac{d}{p}}{1+(2\kappa-\eta)^{+}-\eta}<q\leq\frac{\varepsilon+\frac{d}{p}}{(\varepsilon+\frac{d}{p}-\frac{d}{k})^{+}},$
where we set $\frac{\varepsilon+\frac{d}{p}}{(\varepsilon+\frac{d}{p}-\frac{d}{k})^{+}}=\infty$ if $\varepsilon+\frac{d}{p}-\frac{d}{k}\leq 0$, there exists an increasing function $\beta:[0,\infty)\rightarrow(0,\infty)$ such that the following assertions hold.
  1. $(1)$
    For any $\gamma,\tilde{\gamma}\in\mathscr{P}_{\varepsilon,p*}$ and $),t\in(0,\tau(\gamma)\land\tau(\tilde{\gamma}$)),
[…]
    where, by (2.10),
    (2.12) $\xi(q):=\delta+\frac{dpq-k(d+\varepsilon p)(q-1)}{pqk}\in\big[\eta,1+(2\kappa-\eta)^{+}\big).$
    If $\varepsilon=0$ and $p=\infty$, then for some increasing $\beta:(0,\infty)\rightarrow(0,\infty)$
    (2.13) $\|P_{t}^{*}\gamma-P_{t}^{*}\tilde{\gamma}\|_{\delta,k*}\leq\beta_{t}t^{-\frac{1+\delta}{2}-\frac{d}{2k}}\mathbb{W}_{1}(\gamma,\tilde{\gamma}),\ \ t>0.$
  2. $(2)$
    For any $\theta^{\prime}\in(\theta,\infty),$ $\gamma,\tilde{\gamma}\in\mathscr{P}_{\varepsilon,p*}$ and $),t\in(0,\tau(\gamma)\land\tau(\tilde{\gamma}$)),
    (2.14) $\begin{split}{\rm Ent}(P_{t}^{*}\gamma&|P_{t}^{*}\tilde{\gamma})\leq\beta_{t}(\|\gamma\|_{\varepsilon,p*}+\|\tilde{\gamma}\|_{\varepsilon,p*})^{\frac{2(q-1)}{q}}\\ &\times\bigg(\frac{\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2}}{s_{t}(\theta^{\prime},\gamma)}+\frac{K_{t,\beta}^{(\theta)}(\gamma,\tilde{\gamma})^{2}\mathbb{W}_{q}(\gamma,\tilde{\gamma})^{2}}{[s_{t}(\theta^{\prime},\gamma)\land s_{t}(\theta^{\prime},\tilde{\gamma})]^{([(1+\xi(q))\vee(\delta+\frac{d}{k})]-(2\kappa+1))^{+}}}\bigg).\end{split}$
    In particular, if $\varepsilon=0$ and $p=\infty$, then for some increasing $\beta:(0,\infty)\rightarrow(0,\infty)$
    (2.15) ${\rm Ent}(P_{t}^{*}\gamma|P_{t}^{*}\tilde{\gamma})\leq\frac{\beta_{t}}{t}\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2},\ \ t>0,\ \gamma,\tilde{\gamma}\in\mathscr{P}.$
  3. $(3)$
    Let $(\varepsilon,p)=(\delta,k)$ so that $\xi(q)=\frac{1}{q}(\delta+\frac{d}{k})\in\big[0,1+2\kappa\big)$, the condition (2.9) becomes $\delta<1+2\kappa,$ $\delta+\frac{d}{k}<2+2\kappa$ and (2.10) reduces to $q\in[1,\infty)\cap{(\frac{\delta k+d}{k(1+2\kappa)},1+\frac{d}{\delta k})}$. Then for any $r\in(0,\infty)$ there exists increasing $C_{r}:(0,\infty)\rightarrow(0,\infty)$ such that for any $t>0$ and $\gamma,\tilde{\gamma}\in\hat{\mathscr{P}_{r}}$,
[…]
    ${\rm Ent}(P_{t}^{*}\gamma|P_{t}^{*}\tilde{\gamma})\leq C_{r}(t)\bigg(\frac{\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2}}{t}+\frac{\mathbb{W}_{q}(\gamma,\tilde{\gamma})^{2}}{t^{([(1+\xi(q))\vee(\delta+\frac{d}{k})]-(2\kappa+1))^{+}}}\bigg).$
Remark 2.4.
To see that (2.11) and (2.14) characterize the regularity of the map $\gamma\mapsto P_{t}^{*}\gamma$, let
$P_{t}f(\gamma):=\int_{\mathbb{R}^{d}}f\text{\rm{d}}(P_{t}^{*}\gamma),\ \ \ f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
By Pinsker’s inequality and (3.3) below, there exists a constant $c\in(0,\infty)$ such that
$c\|\gamma-\tilde{\gamma}\|_{var}\leq\|\gamma-\tilde{\gamma}\|_{\delta,k*}\land\sqrt{{\rm Ent}(\gamma|\tilde{\gamma})}$.
So, each of (2.11) and (2.14) implies the local Lipschitz continuity of $\gamma\mapsto P_{t}f(\gamma)$ uniformly in $\|f\|_{\infty}\leq 1$ with respect to $\mathbb{W}_{q}+\mathbb{W}_{2}$:
$\limsup_{\mathscr{P}_{\varepsilon,p*}\ni\tilde{\gamma}\rightarrow\gamma}\sup_{|f|\leq 1}\frac{|P_{t}f(\gamma)-P_{t}f(\tilde{\gamma})|}{\mathbb{W}_{2}(\gamma,\tilde{\gamma})+\mathbb{W}_{q}(\gamma,\tilde{\gamma})}<\infty,\ \ \ \gamma\in\mathscr{P}_{\varepsilon,p*},\ t\in(0,\tau(\gamma)).$
The estimates (2.14)-(2.15) are called entropy-cost inequality or log-Harnack inequality. This type inequalities were first established in [15] for elliptic diffusions on manifolds (possibly with boundary), see [16] for the study of SPDEs, and see [6, 7, 17] for the study of SDEs and McKean-Vlasov SDEs. There are also many other papers concerning log-Harnack inequalities and applications, which we do not mention in details to save space.
We present the following two examples to illustrate Theorem 2.1 and Theorem 2.3, where the kernel $h$ is more singular than the Riesz kernel as considered in previous papers, see [8, 11, 13, 14] and references therein, and covers the Dirac interaction for Nemytskii-type SDEs depending on the density and any order derivatives.
Example 2.5 (Interactions with Arbitrary Singular Order).
Let $b_{t}(x,\mu)=\int_{\mathbb{R}^{d}}h_{t}(x-y)\mu(\text{\rm{d}}y)$ for
$h_{t}(z)=t^{\kappa}h(z),\ \ h(z):=\frac{cz}{|z|^{d+2n_{0}+\varepsilon_{0}}},\ \ 0\neq z\in\mathbb{R}^{d},$
where $\kappa\in[0,\infty)$, $0\neq c\in\mathbb{R}^{d},$ $n_{0}\in\mathbb{Z}_{+}$ and $\varepsilon_{0}\in[0,2)$. When $n_{0}\neq 0$ or $\varepsilon_{0}\geq 1$, $h$ is more singular than Riesz kernels $\bf K$} with $|{\bf K}(z)|\leq c_{0}|z|^{-\beta}$ for some $c_{0}\in(0,\infty)$ and $\beta\in(0,d)$.
Let $\delta=1+2n_{0}$ and $k\in(1,\frac{d}{(d+\varepsilon_{0}-2)^{+}}),$ where $\frac{d}{(d+\varepsilon_{0}-2)^{+}}:=\infty$ if $d+\varepsilon_{0}-2\leq 0$. We claim that assertions in Theorem 2.1 hold for any $\varepsilon\in[0,\delta]$ and $p\in[k,\infty]$ satisfying (2.1), and assertions in Theorem 2.3 hold under (2.9) and $\varepsilon\leq\frac{d(p-1)}{p}.$
In particular, when $\kappa=0$, for any $\mathscr{F}_{0}$-measurable initial value $X_{0}$ with $\mathscr{L}_{X_{0}}\in\hat{\mathscr{P}_{0}}$, the SDE (1.1) has a unique (weak and strong) global solution satisfying
$\sup_{t\in[0,T]}\|\mathscr{L}_{X_{t}}\|_{\delta,k*}<\infty,\ \ \ T\in(0,\infty).$
Proof.
It suffices to verify $h\in\tilde{W}^{-\delta,k}(\mathbb{R}^{d};\mathbb{R}^{d})$ for $\delta=1+2n_{0}$ and $k\in(1,\frac{d}{(d+\varepsilon_{0}-2)^{+}}),$ so that (A) holds.
By $\Delta|\cdot|=\frac{d-1}{|\cdot|}$ and the chain rule, we find a constant $c_{1}\in\mathbb{R}$ such that
(2.16) $h(z)=\nabla\Delta^{n_{0}}W(z),\ \ \ W(z):=c_{1}|z|^{-(d+\varepsilon_{0}-2)}.$
Let $\varphi\in C_{b}^{\infty}(\mathbb{R}^{d})$ such that
$0\leq\varphi\leq 1,\ \ \ \ \varphi|_{B(0,1)}=1,\ \ \ \ \varphi|_{B(0,2)^{c}}=0.$
By (2.16) and $(1-\Delta)^{-s}\nabla=\nabla(1-\Delta)^{-s}$ for $s\in[0,\infty)$, we derive
(2.17) $\big\|(1-\Delta)^{-\frac{1+2n_{0}}{2}}h\big\|_{\tilde{L}^{k}}=\big\|\nabla(1-\Delta)^{-\frac{1}{2}}(1-\Delta)^{-n_{0}}\Delta^{n_{0}}W\big\|_{\tilde{L}^{k}}\leq I_{1}+I_{2},$
where
$I_{1}:=\big\|\nabla(1-\Delta)^{-\frac{1}{2}}(1-\Delta)^{-n_{0}}\Delta^{n_{0}}(\varphi W)\big\|_{L^{k}},$
$I_{2}:=\big\|\nabla(1-\Delta)^{-\frac{1+2n_{0}}{2}}\Delta^{n_{0}}\big((1-\varphi)W\big)\big\|_{\tilde{L}^{k}}.$
Noting that $\nabla(1-\Delta)^{-\frac{1}{2}}$ and $(1-\Delta)^{-n_{0}}\Delta^{n_{0}}$ are bounded in $L^{k}$ for $k\in(1,\infty)$, combining this with
$\|\varphi W\|_{L^{k}}\leq\|1_{B(0,2)}W\|_{L^{k}}<\infty$
for $W(x):=c_{1}|x|^{-(d+\varepsilon_{0}-2)}$ since $\varepsilon_{0}<2$ and $k<\frac{d}{(d+\varepsilon_{0}-2)^{+}}$, we obtain $I_{1}<\infty$.
On the other hand, by the definition of $\varphi$ and $W$, we have
$\big\|P_{t}^{0}\nabla\Delta^{n_{0}}\big((1-\varphi)W\big)\big\|_{L^{\infty}}\leq\big\|\nabla\Delta^{n_{0}}\big((1-\varphi)W\big)\big\|_{L^{\infty}}=<\infty.$
Note that for any $r\in(0,\infty)$ the following formula holds on $\cup_{k\in[1,\infty]}\tilde{W}^{-2r,k}:$
(2.18) $(1-\Delta)^{-r}=\frac{1}{\Gamma(r)}\int_{0}^{\infty}s^{r-1}\text{\rm{e}}^{-s}P_{s}^{0}\text{\rm{d}}s,\ \ \ r>0.$
Then for some constant $c(d,n_{0})\in(0,\infty)$
$I_{2}\leq\omega(d)^{\frac{1}{k}}\|(1-\Delta)^{-\frac{1+2n_{0}}{2}}\nabla\Delta^{n_{0}}\big((1-\varphi)W\big)\|_{L^{\infty}}\leq c(d,n_{0})\int_{0}^{\infty}s^{\frac{1+2n_{0}}{2}-1}\text{\rm{e}}^{-s}<\infty,$
where $\omega(d)$ is the volume of $B(0,1)\subset\mathbb{R}^{d}.$ This together with (2.17) and $I_{1}<\infty$ implies $h\in\tilde{W}^{-\delta,k}(\mathbb{R}^{d};\mathbb{R}^{d})$, and hence finishes the proof. ∎
Next, we consider SDEs whose coefficients depend on any order derivatives of the density function. Let $\mathscr{T}_{0}:=\mathbb{R}$, and for any $i\in\mathbb{N}$ let
$\mathscr{T}_{i}:=\otimes^{i}\mathbb{R}^{d}$
be the space of $i$-tensors over $\mathbb{R}^{d}$. Given $n\in\mathbb{N}$, let
$\mathbb{H}_{n}:=\prod_{i=0}^{n-1}\mathscr{T}_{i},$
which is a finite-dimensional Hilbert space with induced norm $\|\cdot\|_{\mathbb{H}_{n}}.$ For a function $f$ on $\mathbb{R}^{d}$ such that the (n1)(n-1)-th order derivative $\nabla^{n-1}f$ exists and is measurable, where $\nabla^{0}$ denotes the identity map, let
$f^{\langle n}(x):=(\nabla^{i}f(x))_{0\leq i\leq n-1}\in\mathbb{H}_{n},\ \ \ x\in\mathbb{R}^{d}.$
In particular, when $n=1$ we have $f^{\langle 1}=f$.
We consider the following SDE on $\mathbb{R}^{d}$ for a fixed time $T>0$:
(2.19) $\text{\rm{d}}X_{t}=\text{\rm{d}}W_{t}+b_{t}\big(X_{t},\ell_{X_{t}}^{\langle n}(X_{t})\big)\text{\rm{d}}t,\ \ t\in[0,T],$
where $\ell_{X_{t}}(x):=\frac{\mathscr{L}_{X_{t}}(\text{\rm{d}}x)}{\text{\rm{d}}x}$ is the density function of $X_{t},$ and
$b:[0,T]\times\mathbb{R}^{d}\times\mathbb{H}_{n}\rightarrow\mathbb{R}^{d}$
is measurable.
Example 2.6 (Density-Derivative Dependent SDE).
If
(2.20) $\begin{split}&|F_{t}(x,h)-F_{t}(x,\tilde{h})|\leq K_{t}t^{\kappa}\|h-\tilde{h}\|_{\mathbb{H}_{n}},\\ &|F_{t}(x,h)|\leq K_{t}t^{\kappa}(1+\|h\|_{\mathbb{H}_{n}}),\ \ \ t\in[0,\infty),\ h,\tilde{h}\in\mathbb{H}_{n},\ x\in\mathbb{R}^{d}\end{split}$
holds for some $\kappa\in[0,\infty)$ and increasing $K:(0,\infty)\rightarrow(0,\infty).$ Then the density-derivative dependent SDE (2.19) has a unique (weak and strong) global $\mathscr{C}_{\delta,\infty;\delta,\infty}$-solution for any $\delta>d+n-1$ and any initial distribution in $\hat{\mathscr{P}}_{0}$, and
$^{\rm Ent}(P_{t}^{*}\mu|P_{t}^{*}\nu)\leq\frac{c(t,r^{-1})}{t}\mathbb{W}_{2}(\mu,\nu)^{2},\ \ \ t,r>0,\ \mu,\nu\in\hat{\mathscr{P}_{r}}$
holds for some increasing $c:(0,\infty)\times(0,\infty)\rightarrow(0,\infty).$ If moreover $1+2\kappa>d+n-1$, then (2.19) has a unique (weak and strong) global $\mathscr{C}_{0,\infty;\delta,\infty}$-solution for $\delta\in(d+n-1,1+2\kappa)$ and any initial distribution, and when $\frac{1}{2}+\kappa>d+n-1$ there exists increasing $c:(0,\infty)\rightarrow(0,\infty)$ such that
${\rm Ent}(P_{t}^{*}\mu|P_{t}^{*}\nu)\leq\frac{c(t)}{t}\mathbb{W}_{2}(\mu,\nu)^{2},\ \ \ t>0,\ \mu,\nu\in\mathscr{P}.$
Proof.
By Theorem 2.1(3), (1) and Theorem 2.3(2), it suffices to verify (A) for any $\delta>d+n-1$ and $k=\infty,$ since when $1+2\kappa>d+n-1$, we can choose $\delta\in(d+n-1,1+2\kappa)$ so that $\delta=\eta<1+2\kappa$ for $\varepsilon=0$ and $p=\infty$ and when $\kappa>d+n-\frac{3}{2}$ we may choose $\delta\in(d+n-1,\frac{1}{2}+\kappa)$ so that (2.9) holds for $\varepsilon=0$ and $p=\infty$.
For any $0\leq i\leq n-1$, we have
$\nabla^{i}\ell_{X_{t}}=(-1)^{i}(\nabla^{i}{\bf\delta}_{0})*\mathscr{L}_{X_{t}},$
where ${\bf\delta}_{0}$ is the Dirac function. For any $0\leq i\leq n-1$, we find positive constants $\{c_{0}(\delta),c_{i}(\delta)\}\subset(0,\infty)$ such that
$\|\nabla^{i}{\bf\delta}_{0}\|_{\tilde{W}^{-\delta,\infty}}=c_{0}(\delta)\sup_{x\in\mathbb{R}^{d}}\bigg|\int_{0}^{\infty}t^{\frac{\delta}{2}-1}\text{\rm{e}}^{-t}(2\pi t)^{-\frac{d}{2}}\nabla^{i}\{\text{\rm{e}}^{-\frac{|\cdot|^{2}}{2t}}\}(x)\text{\rm{d}}t\bigg|$
$\leq c_{i}(\delta)\int_{0}^{\infty}t^{\frac{\delta-i-d}{2}-1}\text{\rm{e}}^{-t}\text{\rm{d}}t<\infty,\ \ \delta>d+i.$
So, (2.20) implies (A) for $\delta>d+n-1$ and $k=\infty.$
3 Some preparations
We will frequently use the following simple inequality: for any $\alpha_{1},\alpha_{2}\in[0,1)$ and $\alpha\in[0,1-\alpha_{2}],$ there exists $c(\alpha,\alpha_{1},\alpha_{2})\in(0,\infty)$ such that
(3.1) $\int_{0}^{t}s^{-\alpha_{1}}(t-s)^{-\alpha_{2}}\text{\rm{e}}^{-\lambda(t-s)}\text{\rm{d}}s\leq c(\alpha,\alpha_{1},\alpha_{2})t^{1-\alpha-\alpha_{1}-\alpha_{2}}\lambda^{-\alpha},\ \ \ t,\lambda>0.$
Indeed, by the FKG inequality, when $\alpha_{1},\alpha_{2}\in[0,1)$, we have
$\int_{0}^{t}{s^{-\alpha_{1}}}(t-s)^{-\alpha_{2}}\text{\rm{e}}^{-\lambda(t-s)}\text{\rm{d}}s$ $\leq\bigg(\frac{1}{t}\int_{0}^{t}{s^{-\alpha_{1}}}\text{\rm{d}}s\bigg)\int_{0}^{t}(t-s)^{-\alpha_{2}}\text{\rm{e}}^{-\lambda(t-s)}\text{\rm{d}}s$
$=\frac{1}{1-\alpha_{1}}t^{-\alpha_{1}}\int_{0}^{t}s^{-\alpha_{2}}\text{\rm{e}}^{-\lambda s}\text{\rm{d}}s,\ \ t,\lambda>0.$
Then for $\alpha\in[0,1-\alpha_{2})$ the inequality (3.1) follows from Hölder’s inequality
$\int_{0}^{t}s^{-\alpha_{2}}\text{\rm{e}}^{-\lambda s}\text{\rm{d}}s\leq\bigg(\int_{0}^{t}s^{-\frac{\alpha_{2}}{1-\alpha}}\text{\rm{d}}s\bigg)^{1-\alpha}\bigg(\int_{0}^{t}\text{\rm{e}}^{-{\frac{\lambda}{\alpha}}s}\text{\rm{d}}s\bigg)^{\alpha}$
$\leq{\alpha^{\alpha}}\Big(\frac{1-\alpha}{1-\alpha-\alpha_{2}}\Big)^{1-\alpha}t^{{1-\alpha-\alpha_{2}}}\lambda^{-\alpha},\ \ \ t,\lambda>0,$
and when $\alpha=1-\alpha_{2}$, (3.1) is implied by
$\int_{0}^{\infty}s^{-\alpha_{2}}\text{\rm{e}}^{-\lambda s}\text{\rm{d}}s\leq\int_{0}^{\lambda^{-1}}s^{-\alpha_{2}}+\lambda^{\alpha_{2}}\int_{\lambda^{-1}}^{\infty}\text{\rm{e}}^{-\lambda s}\text{\rm{d}}s\leq\frac{2-\alpha_{2}}{1-\alpha_{2}}\lambda^{\alpha_{2}-1},\ \ \lambda>0.$
Lemma 3.1.
Let $P_{t}^{0}$ be the heat semigroup of the Brownian motion. Then there exists increasing $B:[0,\infty)\rightarrow(0,\infty)$ such that
$\|\nabla^{i}P_{t}^{0}\|_{\tilde{W}^{-\delta,k}\rightarrow\tilde{W}^{-\varepsilon,p}}\leq B_{i+\delta-\varepsilon}t^{-\frac{i+\delta-\varepsilon}{2}-\frac{d(p-k)}{2pk}},$
$\ \ t>0,\ 0\leq\varepsilon\leq\delta<\infty,\ 1\leq k\leq p\leq\infty,\ i\in\mathbb{Z}_{+}.$
Proof.
Noting that
$P_{t}^{0}f(x)=\big(2\pi t\big)^{-\frac{d}{2}}\int_{\mathbb{R}^{d}}\text{\rm{e}}^{-\frac{|x-y|^{2}}{2t}}f(y)\text{\rm{d}}y,$
and there exists $L:\mathbb{Z}_{+}\rightarrow(0,\infty)$ such that
$\Big|\nabla^{i}\text{\rm{e}}^{-\frac{|\cdot|^{2}}{2t}}\Big|(x)\leq L_{i}t^{-\frac{i}{2}}\text{\rm{e}}^{-\frac{|x|^{2}}{4t}},\ \ \ t>0,\ i\in\mathbb{Z}_{+},\ x\in\mathbb{R}^{d},$
we find an increasing function $C:\mathbb{Z}_{+}\rightarrow(0,\infty)$ such that
$|(1-\Delta)^{n}\nabla^{i}P_{t}^{0}f|\leq C_{n+i}t^{-\frac{i}{2}-n}P_{2t}^{0}|f|,\ \ t>0,\ i,n\in\mathbb{Z}_{+},f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
On the other hand, by [8, Lemma 5.3], there exists a constant $c_{0}\in(0,\infty)$ such that
$\|P_{2t}^{0}\|_{\tilde{L}^{k}\rightarrow\tilde{L}^{p}}\leq c_{0}t^{-\frac{d(p-k)}{2pk}},\ \ t>0,\ 1\leq k\leq p\leq\infty.$
So,
(3.2) $\|(1-\Delta)^{n}\nabla^{i}P_{t}^{0}\|_{\tilde{L}^{k}\rightarrow\tilde{L}^{p}}\leq C_{n+i}c_{0}t^{-\frac{i}{2}-n-\frac{d(p-k)}{2pk}},\ \ t>0,\ i,n\in\mathbb{Z}_{+}.$
Let $n\in\mathbb{N}$ such that
$\theta_{0}:=n-\frac{\delta-\varepsilon}{2}\in[1,2).$
By (2.18) and (3.2), we find an increasing function $\bar{B}:\mathbb{Z}_{+}\rightarrow(0,\infty)$ such that
$\|\nabla^{i}P_{t}^{0}\|_{\tilde{W}^{-\delta,k}\rightarrow\tilde{W}^{-\varepsilon,p}}=\|(1-\Delta)^{-\frac{\varepsilon}{2}}\nabla^{i}P_{t}^{0}(1-\Delta)^{\frac{\delta}{2}}\|_{\tilde{L}^{k}\rightarrow\tilde{L}^{p}}$
$=\|(1-\Delta)^{n}\nabla^{i}(1-\Delta)^{-\theta_{0}}P_{t}^{0}\|_{\tilde{L}^{k}\rightarrow\tilde{L}^{p}}$
$=\frac{1}{\Gamma(\theta_{0})}\bigg\|\int_{0}^{\infty}s^{\theta_{0}-1}\text{\rm{e}}^{-s}(1-\Delta)^{n}\nabla^{i}P_{t+s}^{0}\text{\rm{d}}s\bigg\|_{\tilde{L}^{k}\rightarrow\tilde{L}^{p}}$
$\leq\frac{C_{n+i}c_{0}}{\Gamma(\theta_{0})}\int_{0}^{\infty}s^{\theta_{0}-1}\text{\rm{e}}^{-s}(t+s)^{-\frac{i}{2}-n-\frac{d(p-k)}{2pk}}\text{\rm{d}}s$
$\leq{\bar{B}_{n+i}}t^{-\frac{i+\delta-\varepsilon}{2}-\frac{d(p-k)}{2pk}},\ \ t>0,\ i\in\mathbb{Z}_{+},\ \delta-\varepsilon\in(2(n-2),2(n-1)],\ n\in\mathbb{N}.$
This implies the desired estimate for some increasing function $B:[0,\infty)\rightarrow(0,\infty).$
Lemma 3.2.
Let $1\leq k\leq p\leq\infty$, $\varepsilon\leq\delta<\infty$, $\lambda\in[0,\infty)$ and $T\in(0,\infty)$.
  1. $(1)$
    The metric space $(\mathscr{P}_{\delta,k*},\ \|\cdot\|_{\delta,k*})$ is complete, and the Borel $\sigma$-field coincides with that induced by the weak topology.
  2. $(2)$
    For any $\gamma\in\mathscr{P}_{\varepsilon,p*},$ the space $(\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T},\rho^{\lambda,T}_{\varepsilon,p;\delta,k})$ is complete, where for $\mathscr{C}_{\varepsilon,p;\delta,k}^{T}$ in (1.4),
    $\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T}:=\big\{\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{T}:\ \mu_{0}=\gamma\big\},$
    $\rho^{\lambda,T}_{\varepsilon,p;\delta,k}(\mu,\nu):=\sup_{t\in(0,T]}\text{\rm{e}}^{-\lambda t}t^{\frac{\delta-\varepsilon}{2}+\frac{d(p-k)}{2pk}}\|\mu_{t}-\nu_{t}\|_{\delta,k*}.$
Proof.
(1) We first observe that the linear space $\tilde{W}^{-\delta,k}$ is Banach under the norm $\|\cdot\|_{\tilde{W}^{-\delta,k}}.$ Noting that
$(1-\Delta)^{-\frac{\delta}{2}}:\ \big(\tilde{W}^{-\delta,k},\|\cdot\|_{\tilde{W}^{-\delta,k}}\big)\rightarrow\big(\tilde{L}^{k},\|\cdot\|_{\tilde{L}^{k}}\big)$
is isometric, it suffices to show that $(\tilde{L}^{k},\|\cdot\|_{\tilde{L}^{k}})$ is complete. Let $\{f_{n}\}_{n\geq 1}\subset\tilde{L}^{k}$ be a Cauchy sequence. Then it is Cauchy sequence in $L^{k}(B(z,1))$ for every $z\in\mathbb{R}^{d}$, so that there exists $f\in L^{k}_{loc}(\mathbb{R}^{d})$ such that
$\lim_{n\rightarrow\infty}\|1_{B(z,1)}(f_{n}-f)\|_{L^{k}}=0,\ \ \ z\in\mathbb{R}^{d}.$
Combining this with $\lim_{n,l\rightarrow\infty}\|f_{n}-f_{l}\|_{\tilde{L}^{k}}=0$, we derive
$\limsup_{n\rightarrow\infty}\|f_{n}-f\|_{\tilde{L}^{k}}=\limsup_{n\rightarrow\infty}\sup_{z\in\mathbb{R}^{d}}\|1_{B(z,1)}(f_{n}-f)\|_{L^{k}}$
[…]
$=\limsup_{n\rightarrow\infty}\limsup_{l\rightarrow\infty}\|f_{n}-f_{l}\|_{\tilde{L}^{k}}=0.$
Hence, $(\tilde{L}^{k},\|\cdot\|_{\tilde{L}^{k}})$ is complete as desired.
Now, we prove the completeness of $(\mathscr{P}_{\delta,k*},\ \|\cdot\|_{\delta,k*})$. Let $\{\mu_{n}\}_{n\geq 1}$ be a Cauchy sequence in $(\mathscr{P}_{\delta,k*},\ \|\cdot\|_{\delta,k*})$. Since $(\mathscr{P}_{\delta,k*},\ \|\cdot\|_{\delta,k*})$ is included by the dual space $(\tilde{W}^{-\delta,k}_{*},\ \|\cdot\|_{\delta,k*})$ of the Banach space $(\tilde{W}^{-\delta,k},\|\cdot\|_{\tilde{W}^{-\delta,k}})$, there exists a unique $\mu\in\tilde{W}^{-\delta,k}_{*}$ such that
$\lim_{n\rightarrow\infty}\|\mu_{n}-\mu\|_{\delta,k*}=\sup_{\|f\|_{\tilde{W}^{-\delta,k}}\leq 1}|\mu_{n}(f)-\mu(f)|=0.$
Since $\mathscr{P}_{\delta,k*}=\tilde{W}_{*}^{-\delta,k}\cap\mathscr{P},$ it remains to show that $\mu\in\mathscr{P}$. When $\delta=0$, we have
$\|f\|_{\tilde{W}^{-0,k}}=\|f\|_{\tilde{L}^{k}}\leq\omega(d)^{\frac{1}{k}}\|f\|_{\infty},$
where $\omega(d)$ is the volume of the unit ball $B(0,1)$. If $\delta>0$, by (2.18) we have
$\|f\|_{\tilde{W}^{-\delta,k}}$ $=\bigg\|\frac{1}{\Gamma(\delta/2)}\int_{0}^{\infty}t^{\frac{\delta}{2}-1}\text{\rm{e}}^{-t}P_{t}^{0}f\text{\rm{d}}t\bigg\|_{\tilde{L}^{k}}$
$\leq\frac{\omega(d)^{\frac{1}{k}}}{\Gamma(\delta/2)}\bigg\|\int_{0}^{\infty}t^{\frac{\delta}{2}-1}\text{\rm{e}}^{-t}P_{t}^{0}f\text{\rm{d}}t\bigg\|_{\tilde{L}^{\infty}}$
$\leq\bigg(\frac{\omega(d)^{\frac{1}{k}}}{\Gamma(\delta/2)}\int_{0}^{\infty}t^{\frac{\delta}{2}-1}\text{\rm{e}}^{-t}\text{\rm{d}}t\bigg)\|f\|_{\infty}.$
In any case, we find a constant $c\in(0,\infty)$ such that $\|\cdot\|_{\tilde{W}^{-\delta,k}}\leq c\|\cdot\|_{\infty}$, hence
(3.3) $\|\mu_{n}-\mu\|_{var}\leq c\|\mu_{n}-\mu\|_{\delta,k*},$
so that
$\lim_{n\rightarrow\infty}\|\mu_{n}-\mu\|_{var}\leq c\lim_{n\rightarrow\infty}\|\mu_{n}-\mu\|_{\delta,k*}=0.$
This together with $\{\mu_{n}\}_{n\geq 1}\subset\mathscr{P}$ implies $\mu\in\mathscr{P}$.
Finally, since $C_{b}(\mathbb{R}^{d})$ is dense in $\tilde{W}^{-\delta,k}$, then for any $f\in\tilde{W}^{-\delta,k}$ and $\mu\in\mathscr{P}_{\delta,k*}$, there exists $\{f_{n}\}_{n\geq 1}\subset C_{b}(\mathbb{R}^{d})$ satisfying $\lim_{n\rightarrow\infty}\|f_{n}-f\|_{\tilde{W}^{-\delta,k}}=0$, which implies
$\lim_{n\rightarrow\infty}|\mu(f_{n})-\mu(f)|\leq\|\mu\|_{\delta,k*}\lim_{n\rightarrow\infty}\|f_{n}-f\|_{\tilde{W}^{-\delta,k}}=0.$
Noting that the Borel $\sigma$-field in $\mathscr{P}_{\delta,k*}$ is induced by
$\big\{\mu\mapsto\mu(f):\ f\in\tilde{W}^{-\delta,k}\big\},$
hence it coincides with the $\sigma$-field induced by the weak topology.
(2) It suffices to prove for $\lambda=0.$ Let $\{\mu^{(n)}\}_{n\geq 1}\subset\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T}$ be a Cauchy sequence with respect to $\rho^{T}_{\varepsilon,p;\delta,k}(\mu,\nu)$. Then for any $t\in(0,T]$, $\{\mu_{t}^{(n)}\}_{n\geq 1}$ is a Cauchy sequence in $\mathscr{P}_{\delta,k*}$, so that by (1), there exists a unique $\mu_{t}\in\mathscr{P}_{\delta,k*}$ such that
(3.4) $\begin{split}&\lim_{n\rightarrow\infty}\rho^{T}_{\varepsilon,p;\delta,k}(\mu^{(n)},\mu)=\lim_{n\rightarrow\infty}\sup_{t\in[0,T]}\lim_{l\rightarrow\infty}t^{\frac{\delta-\varepsilon}{2}+\frac{d(p-k)}{2pk}}\|\mu_{t}^{(n)}-\mu_{t}^{(l)}\|_{\delta,k*}\\ &\leq\lim_{n,l\rightarrow\infty}\rho^{\gamma,T}_{\varepsilon,p;\delta,k}(\mu^{(n)},\mu^{(l)})=0.\end{split}$
It remains to show the weak continuity of $(0,T]\ni t\mapsto\mu_{t}$, which together with (3.4) implies $\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T}.$ For any $f\in C_{b}(\mathbb{R}^{d}),s\in(0,T]$ and $\varepsilon^{\prime}>0$, by (3.3) for $(\varepsilon,p)$ in place of $(\delta,k)$, and applying (3.4), we find large enough $n\geq 1$ such that
$|\mu_{t}^{(n)}(f)-\mu_{t}(f)|\leq c\|f\|_{\infty}\|\mu_{t}^{(n)}-\mu_{t}\|_{\varepsilon,p*}$
$\leq c\|f\|_{\infty}t^{-\frac{\delta-\varepsilon}{2}-\frac{d(p-k)}{2pk}}\rho^{T}_{\varepsilon,p;\delta,k}(\mu^{(n)},\mu)\leq\varepsilon^{\prime},\ \ t\in[s/2,T].$
Combining this with the weak continuity of $(0,T]\ni t\mapsto\mu_{t}^{(n)},$ we derive
$\limsup_{t\rightarrow s}|\mu_{t}(f)-\mu_{s}(f)|\leq\limsup_{t\rightarrow s}\big\{|\mu_{t}^{(n)}(f)-\mu_{s}^{(n)}(f)|+2\varepsilon^{\prime}\big\}=2\varepsilon^{\prime}.$
Since $\varepsilon^{\prime}>0$ is arbitrary, this implies that weak continuity of $(0,T]\ni t\mapsto\mu_{t}.$
For any $T\in(0,\infty)$ and $\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{T}$, consider the SDE
(3.5) $\text{\rm{d}}X_{s,t}^{\mu,x}=b_{t}(X_{s,t}^{\mu,x},\mu_{t})\text{\rm{d}}t+\text{\rm{d}}W_{t},\ \ t\in[s,T],\ X_{s,s}^{\mu,x}=x.$
Lemma 3.3.
Assume (A) and let $\varepsilon\in[0,\delta]$ and $p\in[k,\infty]$ satisfying (2.1). Then for any $T\in(0,\infty)$ and $\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{T}$, the SDE (3.5) is (weakly and strongly) well-posed.
Proof.
By (A) and $\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{T}$, there exists a constant $c\in(0,\infty)$ such that $b_{t}^{\mu}(x):=b_{t}(x,\mu_{t})$ satisfies
$|b_{t}^{\mu}(x)|\leq ct^{\kappa-\frac{\eta}{2}},\ \ t\in(0,T].$
By this and (2.1), we find $q^{\prime}>2$ such that
(3.6) $\|b^{\mu}\|_{\tilde{L}_{q^{\prime}}^{\infty}(T)}=\sup_{z\in\mathbb{R}^{d}}\left(\int_{0}^{T}\|b_{t}^{\mu}1_{B(z,1)}\|_{\infty}^{q^{\prime}}\right)^{\frac{1}{q^{\prime}}}<\infty.$
Then the desired assertion follows from [8, Proposition 5.1]. ∎
For $\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{T}$, under (2.1) we denote
(3.7) $P_{s,t}^{\mu}f(x):=\mathbb{E}[f(X_{s,t}^{\mu,x})],\ \ 0\leq s\leq t\leq T,\ f\in\mathscr{B}_{b}(\mathbb{R}^{d}),x\in\mathbb{R}^{d}.$
The next lemma provides the estimate (1.3) for ${P}_{s,t}^{\mu}$ replacing $P_{t}^{0}$, which is crucial in the proof of the main results.
Lemma 3.4.
[…]
  1. $(1)$
    If $\xi:=\varepsilon_{1}-\varepsilon_{2}+\frac{d(p_{2}-p_{1})}{p_{1}p_{2}}<1\lor\big(2-i+2\kappa-\eta\big)$, then there exists an increasing function $\beta:[0,\infty)\rightarrow(0,\infty)$ such that for any $t\in(0,\infty)$ and $\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{t},$
    (3.8) $\|\nabla^{i}P_{t}^{\mu}\|_{\tilde{W}^{-\varepsilon_{1},p_{1}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}\leq\beta_{t}\exp\big[t\beta_{t}\rho_{\varepsilon,p;\delta,k}^{t}(\mu)^{\theta}\big]t^{-\frac{i+\xi}{2}}.$
  2. $(2)$
    If $\xi:=\varepsilon_{1}-\varepsilon_{2}+\frac{d(p_{2}-p_{1})}{p_{1}p_{2}}<1\lor\big(2-i-(\eta-2\kappa)^{+}\big)$, then there exists an increasing function $\beta:[0,\infty)\rightarrow(0,\infty)$ such that for any $t\in(0,\infty)$ and $\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{t},$
    (3.9) $\|\nabla^{i}P_{s,t}^{\mu}\|_{\tilde{W}^{-\varepsilon_{1},p_{1}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}\leq\beta_{t}\exp\big[t\beta_{t}\rho_{\varepsilon,p;\delta,k}^{t}(\mu)^{\theta}\big](t-s)^{-\frac{i+\xi}{2}},\ \ s\in(0,t).$
Proof.
Let $t\in(0,\infty)$ and $\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{t}.$ We will complete the proof by the following four steps.
(a) We first observe that when $p_{2}>1$,
(3.10) $\sup_{0\leq r\leq s\leq t}(s-r)^{\frac{i}{2}}\|\nabla^{i}{P}_{r,s}^{\mu}\|_{\tilde{W}^{-\varepsilon_{2},p_{2}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}<\infty,\ \ i=0,1.$
By [8, Proposition 5.4], (3.6) implies that for $p_{2}>1$
(3.11) $\sup_{0\leq r\leq s\leq t}(s-r)^{\frac{i}{2}}\|\nabla^{i}P_{r,s}^{\mu}\|_{\tilde{L}^{p_{2}}\rightarrow\tilde{L}^{p_{2}}}<\infty,\ \ i=0,1.$
Moreover, by the Duhamel formula, see [8, Proposition 5.5(2)], we have
(3.12) ${P}_{s,t}^{\mu}f=P_{t-s}^{0}f+\int_{s}^{t}{P}_{s,r}^{\mu}\langle b_{r}(\cdot,\mu_{r}),\nabla P_{t-r}^{0}f\rangle\text{\rm{d}}r,\ \ s\in[0,t],\ f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
We now prove (3.10) by inducing in $l\in\mathbb{N}$ for $\varepsilon_{2}\in[0,lk_{0}],$ where, due to (2.1),
(3.13) $k_{0}:=\frac{1}{2}\land[1-(\eta-2\kappa)^{+}]>0.$
Let $\varepsilon_{2}\in[0,k_{0}].$ By (3.11) and $\|(1-\Delta)^{-\frac{\varepsilon_{2}}{2}}\|_{\tilde{L}^{p_{2}}\rightarrow\tilde{L}^{p_{2}}}<\infty$, we obtain
$c_{1}(\mu,t):=\sup_{0\leq r\leq s\leq t}(s-r)^{\frac{i}{2}}\|\nabla^{i}P_{r,s}^{\mu}\|_{\tilde{L}^{p_{2}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}$
$=\sup_{0\leq r\leq s\leq t}(s-r)^{\frac{i}{2}}\|(1-\Delta)^{-\frac{\varepsilon_{2}}{2}}\nabla^{i}P_{r,s}^{\mu}\|_{\tilde{L}^{p_{2}}\rightarrow\tilde{L}^{p_{2}}}$
$\leq\|(1-\Delta)^{-\frac{\varepsilon_{2}}{2}}\|_{\tilde{L}^{p_{2}}\rightarrow\tilde{L}^{p_{2}}}\sup_{0\leq r\leq s\leq t}(s-r)^{\frac{i}{2}}\|\nabla^{i}P_{r,s}^{\mu}\|_{\tilde{L}^{p_{2}}\rightarrow\tilde{L}^{p_{2}}}<\infty.$
Combining this with (A), (1.3), (3.12) and noting that $s^{\kappa-\frac{\eta}{2}}=s^{-\frac{(\eta-2\kappa)^{+}}{2}}s^{\frac{(2\kappa-\eta)^{+}}{2}}$, we find $c_{2}(\mu,t)\in(0,\infty)$ increasing in $t\in(0,\infty)$, such that
$\|\nabla^{i}{P}_{r,t}^{\mu}\|_{\tilde{W}^{-\varepsilon_{2},p_{2}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}\leq\|\nabla^{i}P_{t-r}^{0}\|_{\tilde{W}^{-\varepsilon_{2},p_{2}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}$
$\quad+K_{t}\int_{r}^{t}s^{\kappa}\|\nabla^{i}{P}_{r,s}^{\mu}\|_{\tilde{L}^{p_{2}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}\|\mu_{s}\|_{\delta,k*}\|\nabla P_{t-s}^{0}\|_{\tilde{W}^{-\varepsilon_{2},p_{2}}\rightarrow\tilde{L}^{p_{2}}}\text{\rm{d}}s$
$\leq{B_{0}}(t-r)^{-\frac{i}{2}}+K_{t}c_{1}(\mu,t)\rho^{t}_{\varepsilon,p;\delta,k}(\mu)\int_{r}^{t}s^{\kappa-\frac{\eta}{2}}(s-r)^{-\frac{i}{2}}(t-s)^{-\frac{\varepsilon_{2}+1}{2}}\text{\rm{d}}s$
$\leq{B_{0}}(t-r)^{-\frac{i}{2}}+c_{2}(\mu,t)\int_{r}^{t}(s-r)^{-\frac{i+(\eta-2\kappa)^{+}}{2}}(t-s)^{-\frac{\varepsilon_{2}+1}{2}}\text{\rm{d}}s.$
By (2.1), (3.13), $\varepsilon_{2}\in[0,k_{0}]$ and $i\leq 1$, we have
$\frac{\varepsilon_{2}+1}{2}\lor\frac{i+(\eta-2\kappa)^{+}}{2}<1,\ \ \ \frac{\varepsilon_{2}+1+(\eta-2\kappa)^{+}}{2}\leq 1,$
so that (3.1) with $\alpha=0$ and $\lambda=0$ implies
$\int_{r}^{t}s^{-\frac{i+(\eta-2\kappa)^{+}}{2}}(t-s)^{-\frac{\varepsilon_{2}+1}{2}}\text{\rm{d}}s\leq c(t-r)^{-\frac{i}{2}},\ \ r\in[0,t)$
for some constant $c\in(0,\infty).$ Therefore, (3.10) holds for $\varepsilon_{2}\in[0,k_{0}].$
Assume that for some $l\in\mathbb{N}$ we have (3.10) for $\varepsilon_{2}\in[0,lk_{0}]$, then for $\varepsilon_{2}\in(lk_{0},(l+1)k_{0}]$,
$c_{l}(\mu,t):=\sup_{0\leq r\leq s\leq t}(s-r)^{\frac{i}{2}}\|\nabla^{i}P_{r,s}^{\mu}\|_{\tilde{W}^{-lk_{0},p_{2}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}$
$=\sup_{0\leq r\leq s\leq t}(s-r)^{\frac{i}{2}}\|(1-\Delta)^{-\frac{\varepsilon_{2}-lk_{0}}{2}}\nabla^{i}P_{r,s}^{\mu}\|_{\tilde{W}^{-lk_{0},p_{2}}\rightarrow\tilde{W}^{-lk_{0},p_{2}}}$
$\leq\|(1-\Delta)^{-\frac{\varepsilon_{2}-lk_{0}}{2}}\|_{\tilde{W}^{-lk_{0},p_{2}}\rightarrow\tilde{\mathbb{W}}^{-lk_{0},p_{2}}}\sup_{0\leq r\leq s\leq t}(s-r)^{\frac{i}{2}}\|\nabla^{i}P_{r,s}^{\mu}\|_{\tilde{W}^{-lk_{0},p_{2}}\rightarrow\tilde{W}^{-lk_{0},p_{2}}}<\infty,$
which is increasing in $t\in(0,\infty)$. By combining this with (A), (1.3), (3.12), we find $c_{l+1}(\mu,t)\in(0,\infty)$ increasing in $t\in(0,\infty)$, such that
$\|\nabla^{i}{P}_{r,t}^{\mu}\|_{\tilde{W}^{-\varepsilon_{2},p_{2}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}\leq\|\nabla^{i}P_{t-r}^{0}\|_{\tilde{W}^{-\varepsilon_{2},p_{2}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}$
$\qquad+K_{t}\int_{r}^{t}s^{\kappa}\|\nabla^{i}{P}_{r,s}^{\mu}\|_{\tilde{W}^{-lk_{0},p_{2}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}\|\mu_{s}\|_{\delta,k*}\|\nabla P_{t-s}^{0}\|_{\tilde{W}^{-\varepsilon_{2},p_{2}}\rightarrow\tilde{W}^{-lk_{0},p_{2}}}\text{\rm{d}}s$
$\leq{B_{0}}t^{-\frac{i}{2}}+K_{t}c_{l}(\mu,t)\rho^{t}_{\varepsilon,p;\delta,k}(\mu){B_{\varepsilon_{2}-lk_{0}}}\int_{r}^{t}s^{\kappa-\frac{\eta}{2}}(s-r)^{-\frac{i}{2}}(t-s)^{-\frac{\varepsilon_{2}-lk_{0}+1}{2}}\text{\rm{d}}s$
$\leq{B_{0}}(t-r)^{-\frac{i}{2}}+c_{l+1}(\mu,t)t^{-\frac{i}{2}},\ \ \ r\in[0,t),\ \varepsilon_{2}\in(lk_{0},(l+1)k_{0}],$
where the last step follows from (3.1) with $\alpha=0$ and $\lambda=0$, since $\frac{\eta+i}{2}-\kappa<1$ for $i\leq 1$, $\frac{\varepsilon_{2}-lk_{0}+1}{2}<1$ for $\varepsilon_{2}-lk_{0}\in[0,k_{0}]$, and $\frac{\varepsilon_{2}-lk_{0}+1}{2}+\frac{(\eta-2\kappa)^{+}}{2}\leq 1.$ Hence, (3.10) holds for all $\varepsilon_{2}\in[0,\infty)$.
(b) We intend to find some increasing function $\beta_{0}:(0,\infty)\rightarrow(0,\infty)$ such that
(3.14) $\begin{split}&h_{r,t}:=\|\nabla^{i}P_{r,t}^{\mu}\|_{\tilde{W}^{-\varepsilon_{2},p_{2}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}\\ &\leq\beta_{0}(t)\exp\big[t\beta_{0}(t)\rho_{\varepsilon,p;\delta,k}^{t}(\mu)^{\theta}\big](t-r)^{-\frac{i}{2}},\ \ \ r\in[0,t),\ i=0,1.\end{split}$
By (1.3), (3.12) and (A), we obtain that
$h_{r,t}\leq\|\nabla^{i}P_{r,t}^{0}\|_{\tilde{W}^{-\varepsilon_{2},p_{2}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}+K_{t}\int_{r}^{t}h_{r,s}s^{\kappa}\|\mu_{s}\|_{\delta,k*}\|\nabla P_{t-s}^{0}\|_{\tilde{W}^{-\varepsilon_{2},p_{2}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}\text{\rm{d}}s$
$\leq{B_{0}}(t-r)^{-\frac{i}{2}}+K_{t}\rho_{\varepsilon,p;\delta,k}^{t}(\mu){B_{0}}t^{(\kappa-\frac{\eta}{2})^{+}}\int_{r}^{t}h_{r,s}(s-r)^{-(\frac{\eta}{2}-\kappa)^{+}}(t-s)^{-\frac{1}{2}}\text{\rm{d}}s,\ \ r\in[0,t).$
Combining this with (3.1) for $\alpha=\frac{1}{\theta}=\frac{1-(\eta-2\kappa)^{+}}{2}$, we find some constant $c(t)\in(0,\infty)$ increasing in $t\in(0,\infty)$, such that
$H_{r,t}(\lambda):=\sup_{s\in(r,t]}(s-r)^{\frac{i}{2}}h_{r,s}\text{\rm{e}}^{-\lambda(s-r)},\ \ \ s\in(r,t]$
satisfies
$H_{r,t}(\lambda)$ $\leq{B_{0}}+K_{t}\rho_{\varepsilon,p;\delta,k}^{t}(\mu){B_{0}}H_{r,t}(\lambda)t^{(\kappa-\frac{\eta}{2})^{+}}$
$\quad\times\sup_{s\in(r,t]}(s-r)^{\frac{i}{2}}\int_{r}^{s}(u-r)^{-\frac{(\eta-2\kappa)^{+}+i}{2}}(s-u)^{-\frac{1}{2}}\text{\rm{e}}^{-\lambda(s-u)}\text{\rm{d}}u$
$\leq{B_{0}}+c(t)\rho_{\varepsilon,p;\delta,k}^{t}(\mu)H_{r,t}(\lambda)\lambda^{-\frac{1}{\theta}},\ \ \ r\in[0,t),\ \lambda>0.$
Taking
$\lambda:={\big(\frac{1}{2}c(t)\rho_{\varepsilon,p;\delta,k}^{t}(\mu)\big)^{\theta},}$
and noting that $H_{r,t}(\lambda)<\infty$ for $p_{2}>1$ due to (3.10), we find some increasing $\beta_{0}:(0,\infty)\rightarrow(0,\infty)$ such that (3.14) holds for any $p_{2}\in(1,\infty]$. Since $\beta_{0}(t)$ is uniformly in $p_{2}>1$, by letting $p_{2}\downarrow 1$ the estimate also holds for $p_{2}=1.$
(c) Let $n\in\mathbb{N}$ such that
$\frac{\xi}{n}<1,\ \ \ \xi:=\varepsilon_{1}-\varepsilon_{2}+\frac{d(p_{2}-p_{1})}{p_{1}p_{2}},$
where we take $n=1$ if $\xi<1$. Below we prove (3.8) by inducing in $0\leq l\leq n$ such that
(3.15) $\|\nabla^{i}P_{t}^{\mu}\|_{\tilde{W}^{-\delta_{l},k_{l}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}\leq\beta_{l}(t)\exp\big[t\beta_{l}(t)\rho_{\varepsilon,p;\delta,k}^{t}(\mu)^{\theta}\big]t^{-\frac{i}{2}-\frac{l\xi}{2n}}.$
for some increasing $\beta_{l}:(0,\infty)\rightarrow(0,\infty)$ and
$\delta_{l}:=\varepsilon_{2}+\frac{l}{n}(\varepsilon_{1}-\varepsilon_{2}),\ \ \ k_{l}:=\frac{np_{1}p_{2}}{(n-l)p_{1}+lp_{2}},\ \ \ 0\leq l\leq n.$
In particular, when $l=n$, (3.15) reduces to the desired inequality (3.8).
By (b), (3.15) holds for $l=0$. Assume that (3.15) holds for some $0\leq l\leq n-1$, where $l=0$ when $\xi<1$, it suffices to verify it for $l+1$ in place of $l$.
To this end, let
$C_{l}(\mu,t):=\beta_{l}(t)\exp\big[t\beta_{l}(t)\rho_{\varepsilon,p;\delta,k}^{t}(\mu)^{\theta}\big]K_{t}B_{\varepsilon_{1}-\varepsilon_{2}}\rho_{\varepsilon,p;\delta,k}^{t}(\mu).$
By (1.3), (3.12), (3.15) and (A), we obtain
$\|\nabla^{i}{P}_{t}^{\mu}\|_{\tilde{W}^{-\delta_{l+1},k_{l+1}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}\leq\|\nabla^{i}P_{t}^{0}\|_{\tilde{W}^{-\delta_{l+1},k_{l+1}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}$
$\qquad+K_{t}\int_{0}^{t}s^{\kappa}\|\nabla^{i}{P}_{s}^{\mu}\|_{\tilde{W}^{-\delta_{l},k_{l}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}\|\mu_{s}\|_{\delta,k*}\|\nabla P_{t-s}^{0}\|_{\tilde{W}^{-\delta_{l+1},k_{l+1}}\rightarrow\tilde{W}^{-\delta_{l},k_{l}}}\text{\rm{d}}s$
$\leq B_{\varepsilon_{1}-\varepsilon_{2}}t^{-\frac{i+(l+1)\xi/n}{2}}+C_{l}(\mu,t)\int_{0}^{t}s^{\kappa-\frac{l\xi}{2n}-\frac{i+\eta}{2}}(t-s)^{-\frac{1+\xi/n}{2}}\text{\rm{d}}s.$
By $\xi<n$, (2.1) and either $\xi<2-i+2\kappa-\eta$ or $\xi<1$ with $l=0$, we have $\frac{1+\xi/n}{2}<1$ and
$\kappa-\frac{l\xi}{2n}-\frac{i+\eta}{2}>-1.$
Moreover, $l+1\leq n$ together with (2.1) implies
$1+\kappa-\frac{l\xi}{2n}-\frac{i+\eta}{2}-\frac{1+\xi/n}{2}\geq-\frac{i+\xi}{2}+1+\kappa-\frac{\eta}{2}>-\frac{i+\xi}{2},$
so by (3.1) with $\alpha=0$, we find $\beta_{l+1}(t)\in(0,\infty)$ which is increasing in $t>0$ such that (3.15) holds for $l+1$ in place of $l$. Hence, (3.8) is proved.
(d) Finally, let $\xi<2-i-(\eta-2\kappa)^{+}$. For the above defined $(\delta_{l},k_{l})_{0\leq l\leq n}$, we intend to find increasing $\beta_{l}:(0,\infty)\rightarrow(0,\infty)$ such that
(3.16) $\|\nabla^{i}P_{r,t}^{\mu}\|_{\tilde{W}^{-\delta_{l},k_{l}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}\leq\beta_{l}(t)\exp\big[t\beta_{l}(t)\rho_{\varepsilon,p;\delta,k}^{t}(\mu)^{\theta}\big](t-r)^{-\frac{i}{2}-\frac{l\xi}{2n}}$
holds for $0\leq l\leq n.$ In particular, when $l=n$ this inequality reduces to the desired (3.9).
By (b), (3.15) holds for $l=0$. Assume that (3.16) holds for some $0\leq l\leq n-1$, it suffices to verify it for $l+1$ in place of $l$.
By (1.3), (3.12), (3.16) and (A), we obtain that for $r\in[0,t)$,
$\|\nabla^{i}{P}_{r,t}^{\mu}\|_{\tilde{W}^{-\delta_{l+1},k_{l+1}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}\leq\|\nabla^{i}P_{t-r}^{0}\|_{\tilde{W}^{-\delta_{l+1},k_{l+1}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}$
$\qquad+K_{t}\int_{r}^{t}s^{\kappa}\|\nabla^{i}{P}_{r,s}^{\mu}\|_{\tilde{W}^{-\delta_{l},k_{l}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}\|\mu_{s}\|_{\delta,k*}\|\nabla P_{t-s}^{0}\|_{\tilde{W}^{-\delta_{l+1},k_{l+1}}\rightarrow\tilde{W}^{-\delta_{l},k_{l}}}\text{\rm{d}}s$
$\leq B_{\varepsilon_{1}-\varepsilon_{2}}(t-r)^{-\frac{i+(l+1)\xi/n}{2}}+{C_{l}(\mu,t)}t^{(\kappa-\frac{\eta}{2})^{+}}\int_{r}^{t}(s-r)^{-(\frac{\eta}{2}-\kappa)^{+}-\frac{i+l\xi/n}{2}}(t-s)^{-\frac{1+\xi/n}{2}}\text{\rm{d}}s.$
Since either $\xi<2-i-(\eta-2\kappa)^{+}$ with $\xi<n$, or $\xi<1$ with $l=0$, we have $\frac{\xi/n+1}{2}<1$ and
$\Big(\frac{\eta}{2}-\kappa\Big)^{+}-\frac{i+l\xi/n}{2}\geq-\Big(\frac{\eta}{2}-\kappa\Big)^{+}-\frac{i+\xi}{2}>-1,$
and (2.1) together with $l+1\leq n$ yields
$1-\frac{i+l\xi/n}{2}-\Big(\frac{\eta}{2}-\kappa\Big)^{+}-\frac{1+\xi/n}{2}>-\frac{i+\xi}{2},$
by (3.1) with $\alpha=0$, we find $\beta_{l+1}(t)\in(0,\infty)$ increasing in $t>0$ such that (3.16) holds for $l+1$ in place $l$. Then the proof is finished. ∎
4 Proof of Theorem 2.1
Let $X_{0}$ be $\mathscr{F}_{0}$-measurable such that $\gamma:=\mathscr{L}_{X_{0}}\in\mathscr{P}_{\varepsilon,p*},$ and let $T\in(0,\infty)$ be fixed. By Lemma 3.3, for any $\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T}$, the SDE
$\text{\rm{d}}X_{t}^{\mu}=b_{t}(X_{t}^{\mu},\mu_{t})\text{\rm{d}}t+\text{\rm{d}}W_{t},\ \ X_{0}^{\mu}=X_{0},\ t\in[0,T]$
has a unique solution. This provides a map
$\Phi:\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T}\rightarrow C^{w}([0,T];\mathscr{P});\ (\Phi_{t}\mu)_{t\in[0,T]}:=(\mathscr{L}_{X_{t}^{\mu}})_{t\in[0,T]}.$
So, the (strong and weak) well-posedness of $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution for (1.1) up to time $T$, if $\Phi$ has a unique fixed point in $\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T}.$
For any $n\in\mathbb{N}$, let $\tau_{n}(\gamma)$ be in (2.4) for some constant $A_{n}\in(0,\infty)$ to be determined and
(4.1) $\tilde{\mathscr{C}}_{\varepsilon,p;\delta,k}^{\gamma,n}:=\Big\{\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,\tau_{n}(\gamma)}:\ \rho_{\varepsilon,p;\delta,k}^{\tau_{n}(\gamma)}(\mu)\leq 2B_{\delta-\varepsilon}{\|\gamma\|_{\varepsilon,p*}}\Big\},$
where $B_{\delta-\varepsilon}\in(0,\infty)$ is in (1.3). Moreover, for any $\lambda\in(0,\infty)$, let
(4.2) $\hat{\mathscr{C}}_{\delta,k}^{\gamma,n,\lambda}:=\Big\{{\mu\in C^{w}([0,n],\mathscr{P})}:\ \mu_{0}=\gamma,\ \rho_{\delta,k}^{n}(\mu):=\sup_{t\in[0,n]}t^{\frac{\delta}{2}+\frac{dp}{2k}}\text{\rm{e}}^{-\lambda t}\|\mu_{t}\|_{\delta,k*}\leq 2{B_{\delta}}\Big\}.$
Lemma 4.1.
Assume (A) and (2.1) with $p>1$. Let $B_{\delta-\varepsilon}\in(0,\infty)$ be in (1.3), and let $\eta:=\delta-\varepsilon+\frac{d(p-k)}{pk}<\kappa+\frac{3}{2}$. Then the following assertions hold.
  1. $(1)$
    For any $T\in(0,\infty)$,
    (4.3) $\Phi:\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T}\rightarrow\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T}.$
  2. $(2)$
    For any $n\in\mathbb{N}$, there exists a constant $A_{n}\in(0,\infty)$ such that
    $\Phi:\tilde{\mathscr{C}}_{\varepsilon,p;\delta,k}^{\gamma,n}\rightarrow\tilde{\mathscr{C}}_{\varepsilon,p;\delta,k}^{\gamma,n}.$
    Moreover, if $\mu$ is a fixed point of $\Phi:\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,\tau_{n}(\gamma)}\rightarrow\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,\tau_{n}(\gamma)},$ then $\mu\in\tilde{\mathscr{C}}_{\varepsilon,p;\delta,k}^{\gamma,n}.$
  3. $(3)$
    If $\eta<\kappa+1$ and (2.1) holds for $\varepsilon=0$ and $p=\infty$, then for any $n\in\mathbb{N}$ there exists a constant $\lambda_{n}\in(0,\infty)$ such that
    $\Phi:\hat{\mathscr{C}}_{\delta,k}^{\gamma,n,\lambda_{n}}\rightarrow\hat{\mathscr{C}}_{\delta,k}^{\gamma,n,\lambda_{n}}.$
    Moreover, if $\mu$ is a fixed point of $\Phi:\mathscr{C}_{0,\infty;\delta,k}^{\gamma,n}\rightarrow\mathscr{C}_{0,\infty;\delta,k}^{\gamma,n},$ then $\mu\in\hat{\mathscr{C}}_{\delta,k}^{\gamma,n,\lambda_{n}}.$
Proof.
(1) Let $T\in(0,\infty)$ and $\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T}.$ We intend to show $\Phi\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T}.$ To this end, let
(4.4) $r=\frac{2}{3+2\kappa},\ \ \ \delta^{\prime}:=r\varepsilon+(1-r)\delta,\ \ \ k^{\prime}:=\frac{pk}{(1-r)p+rk}>1.$
By (2.1) and $\eta<\kappa+\frac{3}{2}$ , we have
(4.5) $\eta<\frac{3+2\kappa}{2}={\frac{2+2\kappa}{2-r}=\frac{1}{r}},$
(4.6) $\delta-\delta^{\prime}+\frac{d(k^{\prime}-k)}{k^{\prime}k}=r\eta,\ \ \ \ \delta^{\prime}-\varepsilon+\frac{d(p-k^{\prime})}{pk^{\prime}}=(1-r)\eta.$
Then $\delta^{\prime}-\varepsilon+\frac{d(p-k^{\prime})}{pk^{\prime}}=(1-r)\eta<2+2\kappa-\eta$, where $k^{\prime}>1$, so that (3.8) holds for $i=0$ due to by Lemma 3.4(1), i.e.
(4.7) $\|{P}_{t}^{\mu}\|_{\tilde{W}^{-\delta^{\prime},k^{\prime}}\rightarrow\tilde{W}^{-\varepsilon,p}}\leq\beta_{t}\exp\big[t\beta_{t}\rho_{\varepsilon,p;\delta,k}^{t}(\mu)^{\theta}\big]t^{-\frac{1-r}{2}\eta},\ \ t\in(0,T],\ \mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T}.$
By (3.12) for $s=0$, for any $\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T}$ we have
$\begin{split}&(\Phi_{t}\mu)(f)=\gamma({P}_{t}^{\mu}f)\\ &=\gamma(P_{t}^{0}f)+\int_{0}^{t}\gamma\big({P}_{s}^{\mu}\langle b_{s}(\cdot,\mu_{s}),\nabla P_{t-s}^{0}f\rangle\big)\text{\rm{d}}s,\ \ t\in[0,T],\ f\in\mathscr{B}_{b}(\mathbb{R}^{d}).\end{split}$
Combining this with (1.3), (4.6), (4.7) and (A), we find a constant $c_{1}\in(0,\infty)$ depending on $T$ and $\mu$ such that
(4.8) $\begin{split}&\|\Phi_{t}\mu\|_{\delta,k*}\leq\|\gamma\|_{\varepsilon,p*}\|P_{t}^{0}\|_{\tilde{W}^{-\delta,k}\rightarrow\tilde{W}^{-\varepsilon,p}}\\ &\qquad+K_{t}\|\gamma\|_{\varepsilon,p*}\int_{0}^{t}s^{\kappa}\|{P}_{s}^{\mu}\|_{\tilde{W}^{-\delta^{\prime},k^{\prime}}\rightarrow\tilde{W}^{-\varepsilon,p}}\|\mu_{s}\|_{\delta,k*}\|\nabla P_{t-s}^{0}\|_{\tilde{W}^{-\delta,k}\rightarrow\tilde{W}^{-\delta^{\prime},k^{\prime}}}\text{\rm{d}}s\\ &\leq B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*}t^{-\frac{\eta}{2}}+c_{1}\int_{0}^{t}s^{\kappa-\frac{2-r}{2}\eta}(t-s)^{-\frac{1}{2}-\frac{r}{2}\eta}\text{\rm{d}}s,\ \ t\in(0,T].\end{split}$
By (4.5), (2.1) and $\eta<\kappa+\frac{3}{2}$ , we have
${\kappa-\frac{2-r}{2}\eta>-1},\ \ -\frac{1}{2}-\frac{r}{2}\eta>-1,\ \ {\kappa-\eta+\frac{1}{2}\geq-\frac{\eta}{2}}.$
By (3.1) for $\alpha=0$ and $\lambda=0$, we find a constant $c\in(0,\infty)$ such that
$\int_{0}^{t}s^{\kappa-\frac{2-r}{2}\eta}(t-s)^{-\frac{1}{2}-\frac{r}{2}\eta}\text{\rm{d}}s\leq ct^{-\frac{\eta}{2}},\ \ t>0.$
Therefore, (4.8) implies $\Phi\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T}.$
(2) Let $\mu\in\tilde{\mathscr{C}}_{\varepsilon,p;\delta,k}^{\gamma,n}$ where the constant $A_{n}$ in $\tau_{n}(\gamma)$ is to be determined such that $\Phi\mu\in\tilde{\mathscr{C}}_{\varepsilon,p;\delta,k}^{\gamma,n}.$ Then
$\begin{split}&K_{n}n^{\frac{(2\kappa-\eta)^{+}}{2}}B_{\delta-\varepsilon}\beta_{n}\rho^{\tau_{n}(\gamma)}_{\varepsilon,p;\delta,k}(\mu)\exp\Big[n\beta_{n}\rho_{\varepsilon,p;\delta,k}^{\tau_{n}(\gamma)}(\mu)^{\theta}\Big]\\ &\leq D_{n}(\gamma):=2K_{n}n^{\frac{(2\kappa-\eta)^{+}}{2}}B_{\delta-\varepsilon}^{2}\|\gamma\|_{\varepsilon,p*}\beta_{n}\exp\Big[n\beta_{n}(2B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*})^{\theta}\Big].\end{split}$
By combining this with (3.12) for $s=0$, (A), (1.3), Lemma 3.4(1), we obtain
(4.9) $\begin{split}&\|\Phi_{t}\mu\|_{\delta,k*}\leq\|\gamma\|_{\varepsilon,p*}\|P_{t}^{0}\|_{\tilde{W}^{-\delta,k}\rightarrow\tilde{W}^{-\varepsilon,p}}\\ &\qquad+{K_{t}}\int_{0}^{t}s^{\kappa}\|\gamma\|_{\varepsilon,p*}\|P_{s}^{\mu}\|_{\tilde{W}^{-\delta^{\prime},k^{\prime}}\rightarrow\tilde{W}^{-\varepsilon,p}}\|\mu_{s}\|_{\delta,k*}\|\nabla P_{t-s}^{0}\|_{\tilde{W}^{-\delta,k}\rightarrow\tilde{W}^{-\delta^{\prime},k^{\prime}}}\text{\rm{d}}s\\ &{\leq B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*}t^{-\frac{\eta}{2}}+\|\gamma\|_{\varepsilon,p*}2K_{n}B_{\delta-\varepsilon}^{2}\|\gamma\|_{\varepsilon,p*}\beta_{n}\exp\Big[n\beta_{n}(2B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*})^{\theta}\Big]}\\ &\qquad\qquad\qquad\qquad\quad\quad\quad{\times\int_{0}^{t}s^{\kappa-\frac{2-r}{2}\eta}(t-s)^{-\frac{1}{2}-\frac{r}{2}\eta}\text{\rm{d}}s}\\ &\leq B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*}t^{-\frac{\eta}{2}}+\|\gamma\|_{\varepsilon,p*}D_{n}(\gamma)ct^{\frac{1}{\theta}-\frac{\eta}{2}},\ \ t\in(0,\tau_{n}(\gamma)],\end{split}$
where $c\in(0,\infty)$ is a constant due to (3.1) for $\alpha=0$ and $\lambda=0$. Taking $A_{n}\in(0,\infty)$ such that
(4.10) $\Big(\frac{B_{\delta-\varepsilon}}{D_{n}(\gamma)c}\Big)^{\theta}\geq\Big(A_{n}\|\gamma\|_{\varepsilon,p*}^{\theta}\text{\rm{e}}^{A_{n}\|\gamma\|_{\varepsilon,p*}^{\theta}}\Big)^{-1},$
by the definition of $\tau_{n}(\gamma)$ we derive
$\|\Phi_{t}\mu\|_{\delta,k*}\leq 2B_{\delta-\varepsilon}{t^{-\frac{\eta}{2}}}\|\gamma\|_{\varepsilon,p*},\ \ \ t\leq\tau_{n}(\gamma).$
So, $\Phi\mu\in\tilde{\mathscr{C}}_{\varepsilon,p;\delta,k}^{\gamma,n}.$
If $\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T}$ is a fixed point of $\Phi$, then $\Phi\mu=\mu$. Noting that $\rho_{\varepsilon,p;\delta,k}^{t}(\mu)$ is non-decreasing in $t$,
$\rho_{\varepsilon,p;\delta,k}^{t+}(\mu):=\lim_{\varepsilon\downarrow 0}\rho_{\varepsilon,p;\delta,k}^{(t+\varepsilon)\land T}(\mu),\ \ \rho_{\varepsilon,p;\delta,k}^{t-}(\mu):=\lim_{\varepsilon\downarrow 0}\rho_{\varepsilon,p;\delta,k}^{(t-\varepsilon)^{+}}(\mu)$
exist and are non-decreasing for $t\in(0,T].$ So, the first inequality in (4.9), (1.3) and Lemma 3.4(1) yield
$\|\mu_{t}\|_{\delta,k*}=\|\Phi_{t}\mu\|_{\delta,k*}\leq B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*}t^{-\frac{\eta}{2}}$
$\quad+{K_{t}B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*}\beta_{t}\exp\big[t\beta_{t}\rho_{\varepsilon,p;\delta,k}^{t-}(\mu)^{\theta}\big]\rho_{\varepsilon,p;\delta,k}^{t-}(\mu)\int_{0}^{t}s^{\kappa-\frac{2-r}{2}\eta}(t-s)^{-\frac{1}{2}-\frac{r}{2}\eta}\text{\rm{d}}s}$
$\leq B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*}t^{-\frac{\eta}{2}}+{K_{t}B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*}t^{\frac{(2\kappa-\eta)^{+}}{2}}\beta_{t}\exp\big[t\beta_{t}\rho_{\varepsilon,p;\delta,k}^{t-}(\mu)^{\theta}\big]\rho_{\varepsilon,p;\delta,k}^{t-}(\mu)ct^{\frac{1}{\theta}-\frac{\eta}{2}}}.$
Thus,
(4.11) $\begin{split}&\rho^{t+}_{\varepsilon,p;\delta,k}(\mu)\leq B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*}\\ &\quad+{K_{t}B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*}t^{\frac{(2\kappa-\eta)^{+}}{2}}\beta_{n}\exp\big[n\beta_{n}\rho_{\varepsilon,p;\delta,k}^{t-}(\mu)^{\theta}\big]\rho_{\varepsilon,p;\delta,k}^{t-}(\mu)ct^{\frac{1}{\theta}}},\ \ t\in(0,\tau_{n}(\gamma)].\end{split}$
Then
$\rho^{0+}_{\varepsilon,p;\delta,k}(\mu)\leq B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*}.$
This and the right continuity of $\rho_{\varepsilon,p;\delta,k}^{t+}(\mu)$ in $t\geq 0$ imply
$s_{0}:=\tau_{n}(\gamma)\land\inf\big\{t\in(0,\tau_{n}(\gamma)]:\ \rho^{t+}_{\varepsilon,p;\delta,k}(\mu)\geq 2B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*}\big\}>0,$
where $\inf\emptyset:=\infty$ by convention. If $s_{0}<\tau_{n}(\gamma)$, by the non-decreasing of $\rho^{t}_{\varepsilon,p;\delta,k}$ in $t\geq 0$, we obtain
$\rho^{s_{0}+}_{\varepsilon,p;\delta,k}(\mu)\geq 2B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*}\geq\rho^{s_{0}-}_{\varepsilon,p;\delta,k}(\mu),$
so that (4.11) yields
$2B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*}\leq\rho^{s_{0}+}_{\varepsilon,p;\delta,k}(\mu)\leq B_{\delta-\varepsilon}\|\gamma\|_{\varepsilon,p*}+D_{n}(\gamma)c\|\gamma\|_{\varepsilon,p*}s_{0}^{\theta}.$
So, (4.10) and the definition of $\tau_{n}(\gamma)$ imply
$s_{0}\geq\Big(\frac{B_{\delta-\varepsilon}}{D_{n}(\gamma)c}\Big)^{\theta}\geq\tau_{n}(\gamma),$
which contradicts to $s_{0}<\tau_{n}(\gamma)$. Hence, $s_{0}\geq\tau_{n}(\gamma)$, which implies $\mu\in\tilde{\mathscr{C}}_{\varepsilon,p;\delta,k}^{\gamma,n}.$
(3) Let $\mu\in\hat{\mathscr{C}}_{\delta,k}^{\gamma,n,\lambda}$, where $\gamma\in\mathscr{P},n\in\mathbb{N}$ and $\lambda\in(0,\infty)$. By (3.12), (1.3), (A) and noting that $\|\gamma\|_{0,\infty*}=1$ for $\gamma\in\mathscr{P}$, we obtain
$\|\Phi_{t}\mu\|_{\delta,k*}$ $\leq B_{\delta}t^{-\frac{\eta}{2}}+K_{n}B_{0}\int_{0}^{t}s^{\kappa}\|\Phi_{s}\mu\|_{\delta,k*}\|\mu_{s}\|_{\delta,k*}(t-s)^{-\frac{1}{2}}\text{\rm{d}}s$
$\leq B_{\delta}t^{-\frac{\eta}{2}}+2B_{0}B_{\delta}K_{n}\int_{0}^{t}s^{\kappa-\frac{\eta}{2}}\|\Phi_{s}\mu\|_{\delta,k*}(t-s)^{-\frac{1}{2}}\text{\rm{d}}s,\ \ t\in(0,n].$
Combining this with (3.1) for $\alpha=\frac{1-(\eta-2\kappa)^{+}}{2}$, when $\eta<1+\kappa$ such that $\kappa-\eta>-1,$ we find a constant $D_{n}\in(0,\infty)$ such that
$H_{n,\lambda}:=\sup_{t\in(0,n]}t^{\frac{\eta}{2}}\|\Phi_{t}\mu\|_{\delta,k*}\text{\rm{e}}^{-\lambda t}$
satisfies
$H_{n,\lambda}$ $\leq B_{\delta}+{2K_{n}B_{0}B_{\delta}}H_{n,\lambda}\sup_{t\in(0,n]}t^{\frac{\eta}{2}}\int_{0}^{t}s^{\kappa-\eta}(t-s)^{-\frac{1}{2}}\text{\rm{e}}^{-\lambda(t-s)}\text{\rm{d}}s$
$\leq B_{\delta}+D_{n}H_{n,\lambda}\lambda^{-\frac{1-(\eta-2\kappa)^{+}}{2}},\ \ \ \lambda>0.$
Noting that (4.3) for $\varepsilon=0$ and $p=\infty$ implies $H_{n,\lambda}<\infty$, taking
$\lambda_{n}:=(2D_{n})^{-\frac{2}{1-(\eta-2\kappa)^{+}}},$
we derive $H_{n,\lambda_{n}}\leq 2B_{\delta-\varepsilon}$, so that $\Phi:\hat{\mathscr{C}}_{\delta,k}^{\gamma,n,\lambda_{n}}\rightarrow\hat{\mathscr{C}}_{\delta,k}^{\gamma,n,\lambda_{n}}.$
If $\mu\in\mathscr{C}_{0,\infty;\delta,k}^{\gamma,n}$ is a fixed point of $\Phi$, then $\mu\in\hat{\mathscr{C}}_{\delta,k}^{\gamma,n,\lambda_{n}}$ can be proved by the same argument as in step (2) to verify that
$s_{0}:=n\land\inf\big\{t\in(0,n]:\ \rho^{t+}_{0,\infty;\delta,k}(\mu)\geq{2B_{\delta}}\big\}=n.$
Let $r$ be in (4.4), by (4.5) and $\eta r<1$, we have
$\frac{1}{\theta_{1}}:=\frac{1}{2}-\frac{r}{2}\eta-\Big(\frac{1-r}{2}\eta-\kappa\Big)^{+}=\Big(\frac{1}{2}-\frac{r}{2}\eta\Big)\land\frac{1}{\theta}>0.$
We have the following estimate on the Lipschitz continuity of $\Phi$ under $\rho_{\varepsilon,p;\delta,k}^{\lambda,T}.$
Lemma 4.2.
Assume (A) and (2.1) with $p>1$. Then there exists increasing $C:(0,\infty)\rightarrow(0,\infty)$ such that
$\rho_{\varepsilon,p;\delta,k}^{\lambda,T}(\Phi\mu,\Phi\nu)\leq\lambda^{-\frac{1}{\theta_{1}}}\|\gamma\|_{\varepsilon,p*}C(T)\exp\Big[C(T)\big(\rho_{\varepsilon,p;\delta,k}^{T}(\mu)+\rho_{\varepsilon,p;\delta,k}^{T}(\nu)\big)^{\theta}\Big]\rho_{\varepsilon,p;\delta,k}^{\lambda,T}(\mu,\nu),$
$\qquad\ T\in(0,\infty),\ \mu,\nu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T},\ \lambda\geq\big(C(T)\rho_{\varepsilon,p;\delta,k}^{T}(\mu)\big)^{\theta}.$
Proof.
Let $r,\delta^{\prime}$ and $k^{\prime}$ be in (4.4). By (3.12) for $s=0$, (A), (1.3) and (4.7), we obtain
$\|P_{t}^{\mu}-P_{t}^{\nu}\|_{\tilde{W}^{-\delta^{\prime},k^{\prime}}\rightarrow\tilde{W}^{-\varepsilon,p}}=\sup_{f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\|f\|_{\tilde{W}^{-\delta^{\prime},k^{\prime}}}\leq 1}\|P_{t}^{\mu}f-P_{t}^{\nu}f\|_{\tilde{W}^{-\varepsilon,p}}$
$\leq\int_{0}^{t}\sup_{f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\|f\|_{\tilde{W}^{-\delta^{\prime},k^{\prime}}}\leq 1}\Big[\big\|(P_{s}^{\mu}-P_{s}^{\nu})\langle b_{s}(\cdot,\mu_{s}),\nabla P_{t-s}^{0}f\rangle\big\|_{\tilde{W}^{-\varepsilon,p}}$
$\qquad\qquad\qquad\qquad\qquad\qquad+\big\|P_{s}^{\nu}\langle b_{s}(\cdot,\mu_{s})-b_{s}(\cdot,\nu_{s}),\nabla P_{t-s}^{0}f\rangle\big\|_{\tilde{W}^{-\varepsilon,p}}\Big]\text{\rm{d}}s$
$\leq B_{0}K_{t}\rho_{\varepsilon,p;\delta,k}^{t}(\mu)\int_{0}^{t}s^{\kappa-\frac{\eta}{2}}\|P_{s}^{\mu}-P_{s}^{\nu}\|_{\tilde{W}^{-\delta^{\prime},k^{\prime}}\rightarrow\tilde{W}^{-\varepsilon,p}}(t-s)^{-\frac{1}{2}}\text{\rm{d}}s$
$\quad+B_{0}K_{t}c_{0}(t)\exp\big[c_{0}(t)\rho_{\varepsilon,p;\delta,k}^{t}(\nu)^{\theta}\big]\rho_{\varepsilon,p;\delta,k}^{\lambda,t}(\mu,\nu)\int_{0}^{t}s^{\kappa-\frac{2-r}{2}\eta}\text{\rm{e}}^{\lambda s}(t-s)^{-\frac{1}{2}}\text{\rm{d}}s.$
Combining this with (3.1) for $\alpha=\frac{1}{\theta}$ and $\alpha=0$ respectively, $\alpha_{1}=(\frac{2-r}{2}\eta-\kappa)^{+}<1$ due to (4.5), and $\alpha_{2}=\frac{1}{2}$, we find $c_{1}(t)\in(0,\infty)$ increasing in $t>0$ such that
$H_{t}(\lambda):=\sup_{u\in(0,t]}\text{\rm{e}}^{-\lambda u}u^{\frac{1-r}{2}\eta}\|P_{u}^{\mu}-P_{u}^{\nu}\|_{\tilde{W}^{-\delta^{\prime},k^{\prime}}\rightarrow\tilde{W}^{-\varepsilon,p}}$
$\leq B_{0}K_{t}\Big[\rho_{\varepsilon,p;\delta,k}^{t}(\mu)H_{t}(\lambda)+c_{0}(t)\exp\big[c_{0}(t)\rho_{\varepsilon,p;\delta,k}^{t}(\nu)^{\theta}\big]\rho_{\varepsilon,p;\delta,k}^{\lambda,t}(\mu,\nu)\Big]$
$\qquad\times\sup_{u\in(0,t]}u^{\frac{1-r}{2}\eta+(\kappa-\frac{2-r}{2}\eta)^{+}}\int_{0}^{u}s^{-(\frac{2-r}{2}\eta-\kappa)^{+}}(u-s)^{-\frac{1}{2}}\text{\rm{e}}^{-\lambda(u-s)}\text{\rm{d}}s$
$\leq c_{1}(t)\rho_{\varepsilon,p;\delta,k}^{t}(\mu)H_{t}(\lambda)\lambda^{-\frac{1}{\theta}}+c_{1}(t)\exp\big[c_{0}(t)\rho_{\varepsilon,p;\delta,k}^{t}(\nu)^{\theta}\big]\rho_{\varepsilon,p;\delta,k}^{\lambda,t}(\mu,\nu)$
holds for all $\lambda>0$. By (3.8) for $i=0$, $(\varepsilon_{1},p_{1})=(\delta^{\prime},k^{\prime})$ and $(\varepsilon_{2},p_{2})=(\varepsilon,p)$, we have $H_{T}(\lambda)<\infty$ for $\lambda>0$. So, taking
(4.12) $\bar{\lambda}(T,\mu)=\big(2c_{1}(T)\rho_{\varepsilon,p;\delta,k}^{T}(\mu)\big)^{\theta},$
we find a constant $c_{2}(t)\in(0,\infty)$ increasing in $t$ such that for any $\lambda\geq\bar{\lambda}(T,\mu)$,
(4.13) $\begin{split}H_{t}(\lambda)&\leq c_{2}(t)\exp\Big[c_{2}(t)\big(\rho_{\varepsilon,p;\delta,k}^{t}(\nu)+\rho_{\varepsilon,p;\delta,k}^{t}(\mu)\big)^{\theta}\Big]\rho_{\varepsilon,p;\delta,k}^{\lambda,t}(\mu,\nu),\\ &\qquad\qquad\ \ \ t\in(0,T],\ \mu,\nu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,t}.\end{split}$
Let $(r,\delta^{\prime},k^{\prime})$ be in (4.4). By (A) and (3.12) for $s=0$, we have
$\|\Phi_{t}\mu-\Phi_{t}\nu\|_{\delta,k*}=\sup_{f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\|f\|_{\tilde{W}^{-\delta,k}}\leq 1}\big|\gamma(P_{t}^{\mu}-P_{t}^{\nu})f\big|$
$\leq\|\gamma\|_{\varepsilon,p*}\int_{0}^{t}\sup_{f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\|f\|_{\tilde{W}^{-\delta,k}}\leq 1}\big\|P_{t}^{\mu}\langle b_{s}(\cdot,\mu_{s}),\nabla P_{t-s}^{0}f\rangle-P_{t}^{\nu}\langle b_{s}(\cdot,\nu_{s}),\nabla P_{t-s}^{0}f\rangle\big\|_{\tilde{W}^{-\varepsilon,p}}\text{\rm{d}}s$
$\leq\|\gamma\|_{\varepsilon,p*}K_{t}\int_{0}^{t}s^{\kappa}\|P_{s}^{\mu}-P_{s}^{\nu}\|_{\tilde{W}^{-\delta^{\prime},k^{\prime}}\rightarrow\tilde{W}^{-\varepsilon,p}}\|\mu_{s}\|_{\delta,k*}\|\nabla P_{t-s}^{0}\|_{\tilde{W}^{-\delta,k}\rightarrow\tilde{W}^{-\delta^{\prime},k^{\prime}}}\text{\rm{d}}s$
$\quad+\|\gamma\|_{\varepsilon,p*}K_{t}\int_{0}^{t}s^{\kappa}\|P_{s}^{\nu}\|_{\tilde{W}^{-\delta^{\prime},k^{\prime}}\rightarrow\tilde{W}^{-\varepsilon,p}}\|\mu_{s}-\nu_{s}\|_{\delta,k*}\|\nabla P_{t-s}^{0}\|_{\tilde{W}^{-\delta,k}\rightarrow\tilde{W}^{-\delta^{\prime},k^{\prime}}}\text{\rm{d}}s.$
Combining this with (1.3), (4.6) and (4.7) we find $c_{3}(t)\in(0,\infty)$ increasing in $t$ such that
$\begin{split}&\|\Phi_{t}\mu-\Phi_{t}\nu\|_{\delta,k*}\\ &\leq c_{3}(t)\|\gamma\|_{\varepsilon,p*}\rho_{\varepsilon,p;\delta,k}^{t}(\mu)\int_{0}^{t}s^{\kappa-\frac{\eta}{2}}\|P_{s}^{\mu}-P_{s}^{\nu}\|_{\tilde{W}^{-\delta^{\prime},k^{\prime}}\rightarrow\tilde{W}^{-\varepsilon,p}}(t-s)^{-\frac{1}{2}-\frac{r}{2}\eta}\text{\rm{d}}s\\ &+c_{3}(t)\|\gamma\|_{\varepsilon,p*}\exp\big[c_{0}(t)\rho_{\varepsilon,p;\delta,k}^{t}(\mu)^{\theta}\big]\int_{0}^{t}\|\mu_{s}-\nu_{s}\|_{\delta,k*}s^{\kappa-\frac{1-r}{2}\eta}(t-s)^{-\frac{1}{2}-\frac{r}{2}\eta}\text{\rm{d}}s.\end{split}$
Thus, by (4.13), and noting that $1+2\kappa-\eta\geq 0$ due to (2.1), we find a constant $c_{4}(t)\in(0,\infty)$ increasing in $t$ such that
(4.14) $\begin{split}&\rho_{\varepsilon,p;\delta,k}^{\lambda,T}(\Phi\mu,\Phi\nu)=\sup_{t\in(0,T]}t^{\frac{\eta}{2}}\text{\rm{e}}^{-\lambda t}\|\Phi_{t}\mu-\Phi_{t}\nu\|_{\delta,k*}\\ &\leq\|\gamma\|_{\varepsilon,p*}c_{3}(T)\exp\Big[c_{4}(T)\big(\rho_{\varepsilon,p;\delta,k}^{T}(\nu)+\rho_{\varepsilon,p;\delta,k}^{T}(\mu)\big)^{\theta}\Big]\rho_{\varepsilon,p;\delta,k}^{\lambda,T}(\mu,\nu)\\ &\qquad\times\sup_{t\in(0,T]}{t^{\frac{\eta}{2}}}t^{(\kappa-\frac{2-r}{2}\eta)^{+}}\int_{0}^{t}s^{-(\frac{2-r}{2}\eta-\kappa)^{+}}(t-s)^{-\frac{1}{2}-\frac{r}{2}\eta}\text{\rm{e}}^{-\lambda(t-s)}\text{\rm{d}}s,\\ &{\qquad\qquad\ \ \ t\in(0,T],\ \mu,\nu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T},\ \lambda\geq\bar{\lambda}(T,\mu).}\end{split}$
By (3.1) for $\alpha=\theta_{1}$, $\alpha_{1}=(\frac{2-r}{2}\eta-\kappa)^{+}<1$ and $\alpha_{2}=\frac{1}{2}+\frac{r}{2}\eta<1$ due to (4.5), we find a constant $c_{5}(T)\in(0,\infty)$ increasing in $T$ such that
$\sup_{t\in(0,T]}t^{\frac{\eta}{2}+(\kappa-\frac{2-r}{2}\eta)^{+}}\int_{0}^{t}s^{-(\frac{2-r}{2}\eta-\kappa)^{+}}(t-s)^{-\frac{1}{2}-\frac{r}{2}\eta}\text{\rm{e}}^{-\lambda(t-s)}\text{\rm{d}}s$
${\leq c_{5}(T)\lambda^{-\theta_{1}},\ \ \ \lambda\geq\bar{\lambda}(T,\mu),\ t\in(0,T].}$
Combining this with (4.12) and (4.14), we derive the desired estimate for some $C(T)\in(0,\infty)$ increasing in $T$. ∎
We are now ready to prove the following result.
Proposition 4.3.
Assume (A) and let $\varepsilon\in[0,\delta]$ and $p\in[k,\infty]$ satisfying (2.1).
  1. $(1)$
    For any $\mathscr{F}_{0}$-measurable initial value $X_{0}$ with $\gamma=\mathscr{L}_{X_{0}}\in\mathscr{P}_{\varepsilon,p*}$, (1.1) has a unique maximal (weak and strong) $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution.
  2. $(2)$
    For any $n\in\mathbb{N}$, there exist $A_{n},\lambda_{n}\in(0,\infty)$ such that (2.4) and (2.5) hold.
  3. $(3)$
    There exists an increasing function $C_{\gamma}:[1,\infty)\rightarrow(0,\infty)$ which does not depend on $\gamma$ when $\varepsilon=0$ and $p=\infty$, such that (2.3) holds.
Proof.
According to the proof of [8, Theorem 2.1], the first assrtion follows from the second. So, it suffices to show that for any $n\in\mathbb{N}$ and initial value $X_{0}$ with $\gamma\in\mathscr{L}_{X_{0}}\in\mathscr{P}_{\varepsilon,p*},$ (1.1) has a unique weak/strong $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution up to $\tau_{n}(\gamma)$ such that
(4.15) $\mathbb{E}\bigg[\sup_{t\in(0,\tau_{n}(\gamma)]}|X_{t}|^{q}\bigg|\mathscr{F}_{0}\bigg]\leq c_{n,q}(\gamma)\big(1+|X_{0}|^{q}\big),\ \ q\in(0,\infty),$
where $c_{n,q}(\gamma)\in(0,\infty)$, which does not depend on $\gamma$ when $\varepsilon=0$ and $p=\infty$. Below, we simply denote $\tau_{n}=\tau_{n}(\gamma).$
(a) By Lemma 3.3 and (4.3), (1.1) has a unique weak/strong $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution up to $\tau_{n}$ provided $\Phi:\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,\tau_{n}}\rightarrow\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,\tau_{n}}$ has a unique fixed point.
By Lemma 4.1, all fixed points in $\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,\tau_{n}}$ of $\Phi$ are included in $\tilde{\mathscr{C}}_{\varepsilon,p;\delta,k}^{\gamma,n}$ when $\varepsilon>0$ or $p<\infty$, and in $\hat{\mathscr{C}}_{\delta,k}^{\gamma,n,\lambda_{n}}$ when $\varepsilon=0$ and $p=\infty$. So, (2.5) holds for any (weak) $\mathscr{C}_{\varepsilon,p;\delta,k}$-solutions of (1.1) with initial distribution $\gamma$ up to time $\tau_{n}$. By Lemma 3.2 and the contractive fixed point theorem, it suffices to find $\lambda\in(0,\infty)$ such that the map $\Phi$ is contractive under the metric $\rho^{\lambda,\tau_{n}}_{\varepsilon,p;\delta,k}$ on $\tilde{\mathscr{C}}_{\varepsilon,p;\delta,k}^{\gamma,n}$ when $\varepsilon>0$ or $p<\infty$, and on $\hat{\mathscr{C}}_{\delta,k}^{\gamma,n,\lambda_{n}}$ otherwise.
Let $\varepsilon>0$ or $p<\infty$. By Lemma 4.2 and (4.1), we can find $C_{n,\gamma}\in(0,\infty)$ such that
$\rho_{\varepsilon,p;\delta,k}^{\lambda,\tau_{n}}(\Phi\mu,\Phi\nu)\leq\lambda^{-\frac{1}{\theta_{1}}}C_{n,\gamma}\rho_{\varepsilon,p;\delta,k}^{\lambda,\tau_{n}}(\mu,\nu),\ \ \mu,\nu\in\tilde{\mathscr{C}}_{\varepsilon,p;\delta,k}^{\gamma,n},\ \ \lambda\geq C_{n,\gamma}.$
Similarly, when $\varepsilon=0$ and $p=\infty$, by Lemma 4.2 and (4.2), we find a constant $C_{n}\in(0,\infty)$ uniformly in $\gamma\in\mathscr{P}$ such that
$\rho_{\varepsilon,p;\delta,k}^{\lambda,n}(\Phi\mu,\Phi\nu)\leq\lambda^{-\frac{1}{\theta_{1}}}C_{n}\rho_{\varepsilon,p;\delta,k}^{\lambda,n}(\mu,\nu),\ \ \mu,\nu\in\hat{\mathscr{C}}_{\delta,k}^{\gamma,n,\lambda_{n}},\lambda\geq C_{n}.$
So, in any case $\Phi$ is contractive under the metric $\rho^{\lambda,\tau_{n}}_{\varepsilon,p;\delta,k}$ when $\lambda>0$ is large enough. So, (1.1) has a unique weak/strong $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution up to $\tau_{n}$.
(b) Let $X_{t}$ be the unique solution up to time $\tau_{n}$. Then
$(\mu_{t}=\mathscr{L}_{X_{t}})_{t\in[0,\tau_{n}]}\in\begin{cases}\tilde{\mathscr{C}}_{\varepsilon,p;\delta,k}^{\gamma,n},&\text{if}\ \varepsilon>0\ \text{or}\ p<\infty,\\ \hat{\mathscr{C}}_{\delta,k}^{\gamma,n,\lambda_{n}},&\text{if}\ \varepsilon=0,p=\infty.\end{cases}$
Then by (2.1) and (2.5), there exists $q^{\prime}>2$ such that $\|b^{\mu}\|_{\tilde{L}_{q^{\prime}}^{\infty}(\tau_{n})}$ in (3.6) is bounded above by a constant $D_{n,\gamma}\in(0,\infty)$ depending on $n$ and $\|\gamma\|_{\varepsilon,p*}$. So, (4.15) follows from [17, Theorem 1.3.1], where $c_{n,q}(\gamma)$ is uniform in $\gamma\in\mathscr{P}$ when $\varepsilon=0$ and $p=\infty$, since $\|\gamma\|_{0,\infty*}=1$. ∎
Proof of Theorem 2.1.
By Proposition 4.3, it remains to prove Theorem 2.1(3). Let $\eta=0$ and $\gamma=\gamma_{0}*N_{0,r}$ for some $r>0$ and $\gamma_{0}\in\mathscr{P}$. As explained above, we only need to prove that for any $T\in(0,\infty)$,
$\Phi:\mathscr{C}_{\delta,k;\delta,k}^{\gamma,T}\rightarrow\mathscr{C}_{\delta,k;\delta,k}^{\gamma,T}$
has a unique fixed point. To this end, we consider the shifted drift
$\hat{b}_{t}^{(r)}(x,\mu):=1_{[r,T]}(t){b_{(t-r)^{+}}(x,\mu)},\ \ \ t\in[0,T+r],\ x\in\mathbb{R}^{d},\ \mu\in\mathscr{P}_{\delta,k*}.$
So, for any constant $\kappa^{\prime}>\delta+\frac{d}{k}$, (A) holds for $(\hat{b}^{(r)},\kappa^{\prime})$ in place of $(b,\kappa)$. Then by Proposition 4.3 for $\varepsilon=0$ and $p=\infty$, for any $\mathscr{F}_{0}$-measurable initial value $\hat{X}_{0}$ with $\mathscr{L}_{\hat{X}_{0}}=\gamma_{0}$, the SDE
(4.16) $\text{\rm{d}}\hat{X}_{t}=\hat{b}_{t}^{(r)}(\hat{X}_{t},\mathscr{L}_{\hat{X}_{t}})\text{\rm{d}}t+\text{\rm{d}}W_{t},\ \ t\in[0,T+r]$
has a unique solution such that
(4.17) $\sup_{t\in(0,T+r]}\sup_{\gamma_{0}\in\mathscr{P}}t^{\frac{\delta}{2}+\frac{d}{2k}}\|\mathscr{L}_{\hat{X}_{t}}\|_{\delta,k*}<\infty.$
So, $(\mu_{t}:=\mathscr{L}_{\hat{X}_{r+t}})_{t\in[0,T]}$ is a fixed point in $\mathscr{C}_{\delta,k;\delta,k}^{\gamma,T}.$
On the other hand, if $\nu\in\mathscr{C}_{\delta,k;\delta,k}^{\gamma,T}$ is a fixed point of $\Phi$, by Lemma 3.3, the SDE
$\text{\rm{d}}X_{t}^{\nu}=b_{t}(X_{t}^{\nu},\nu_{t})\text{\rm{d}}t+\text{\rm{d}}W_{t+r},\ \ X_{0}^{\nu}=\hat{X}_{0}+W_{r},\ t\in[0,T]$
has a a unique solution, and $\mathscr{L}_{X_{t}^{\nu}}=\nu_{t}$ since $\nu$ is a fixed point of $\Phi$. Thus,
$\tilde{X}_{t}:=1_{[0,r]}(t)(\hat{X}_{0}+W_{t})+1_{(r,T+r]}(t)X_{t-r}^{\nu},\ \ t\in[0,T+r]$
solves (4.16). So, the uniqueness of (4.16) implies $\nu_{t}=\mathscr{L}_{\tilde{X}_{r+t}}=\mathscr{L}_{\hat{X}_{r+t}}=\mu_{t},\ \ t\in[0,T]$. Noting that $P_{t}^{*}\gamma=\mathscr{L}_{\hat{X}_{r+t}}$ for $t\geq 0$, (4.17) implies (2.6). ∎
5 Proof of Theorem 2.3
According to Theorem 2.1(3), Theorem 2.3(3) follows from Theorem 2.3(1)-(2). So, we only prove the first two assertions in Theorem 2.3.
By Theorem 2.1, for any $\gamma\in\mathscr{P}_{\varepsilon,p*}$ and $T\in(0,\tau(\gamma))$, we have $(P_{t}^{*}\gamma)_{t\in[0,T]}\in\mathscr{C}_{\varepsilon,p;\delta,k}^{\gamma,T}.$ Let $P_{s,t}^{\gamma}={P}_{s,t}^{\mu}$ be defined in (3.7) for $\mu_{t}=P_{t}^{\ast}\gamma$, i.e.
$P_{s,t}^{\gamma}f(x)=\mathbb{E}[f({X}_{s,t}^{\gamma,x})],\ \ 0\leq s\leq t<\tau(\gamma),\ f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\ x\in\mathbb{R}^{d},$
where for fixed $(s,x)\in[0,\tau(\gamma))\times\mathbb{R}^{d}$, $({X}_{s,t}^{\gamma,x})_{t\in[s,\tau(\gamma))}$ is the unique solution to the SDE
$\text{\rm{d}}{X}_{s,t}^{\gamma,x}=b_{t}({X}_{s,t}^{\gamma,x},P_{t}^{\ast}\gamma)\text{\rm{d}}t+\text{\rm{d}}W_{t},\ \ {X}_{s,s}^{\gamma,x}=x,\ t\in[s,\tau(\gamma)).$
Simply denote $P_{t}^{\gamma}=P_{0,t}^{\gamma}$ for $t\in[0,\tau(\gamma))$.
We now prove that (2.9) implies
(5.1) $1+(2\kappa-\eta)^{+}-\eta>\Big(\varepsilon+\frac{d}{p}-\frac{d}{k}\Big)^{+},$
which is claimed in Theorem 2.3.
If $\eta<2\kappa$, then $\eta<1\lor(\frac{1}{2}+\kappa)$ implies
$1+(2\kappa-\eta)^{+}-\eta=1+2\kappa-2\eta>0,$
and by $\delta<1+(2\kappa-\eta)^{+}=1+2\kappa-\eta$,
$1+(2\kappa-\eta)^{+}-\eta-\Big(\varepsilon+\frac{d}{p}-\frac{d}{k}\Big)=1+2\kappa-2\eta+\eta-\delta=1+2\kappa-\eta-\delta>0,$
so that (5.1) holds.
If $\eta\geq 2\kappa$ then (2.9) implies $\kappa<\frac{1}{2}$ and $\delta<1$, so that
$1+(2\kappa-\eta)^{+}-\eta=1-\eta>0,$
and
$1+(2\kappa-\eta)^{+}-\eta-\Big(\varepsilon+\frac{d}{p}-\frac{d}{k}\Big)=1-\eta+\eta-\delta=1-\delta>0.$
Hence, (2.9) implies (5.1).
Proof of Theorem 2.3(1).
All constants $\{c_{i}(t)\}_{i\geq 1}\subset(0,\infty)$ below are increasing in $t>0$. For any $\gamma\in\mathscr{P}_{\varepsilon,p*},$ define $P_{t}^{\gamma*}:\mathscr{P}\rightarrow\mathscr{P}$ by
$(P_{t}^{\gamma\ast}\nu)(f):=\int_{\mathbb{R}^{d}}\big(P_{t}^{\gamma}f(x)\big)\nu(\text{\rm{d}}x),\ \ f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\ t\in[0,\tau(\gamma)),\ \nu\in\mathscr{P}.$
Then $P_{t}^{*}\gamma=P_{t}^{\gamma*}\gamma$ for $\gamma\in\mathscr{P}_{\varepsilon,p*},$ so that
(5.2) $\|P_{t}^{*}\gamma-P_{t}^{*}\tilde{\gamma}\|_{\delta,k*}\leq\|P_{t}^{\tilde{\gamma}*}\gamma-P_{t}^{\tilde{\gamma}*}\tilde{\gamma}\|_{\delta,k*}+{\|P_{t}^{\gamma*}\gamma-P_{t}^{\tilde{\gamma}*}\gamma\|_{\delta,k*}}.$
In the following, we estimate these two terms respectively for fixed $\gamma,\tilde{\gamma}\in\mathscr{P}_{\varepsilon,p*}$.
[…]
(5.3) $\mathbb{W}_{q}(\gamma,\tilde{\gamma})=\bigg(\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{q}\pi(\text{\rm{d}}x,\text{\rm{d}}y)\bigg)^{\frac{1}{q}}.$
By $\varepsilon\in[0,\frac{d(p-1)}{p}]$, we have $p_{0}:=\frac{dp}{d+\varepsilon p}\in[1,p]$ and
$\frac{1}{p_{0}}=\frac{1}{p}+\frac{\varepsilon}{d}.$
Then the Sobolev embedding theorem implies
$\|\cdot\|_{\tilde{W}^{-\varepsilon,p}}\leq c_{0}\|\cdot\|_{\tilde{L}^{p_{0}}}$
for some constant $c_{0}\in(0,\infty),$ so that
(5.4) $\|\gamma\|_{p_{0}*}:=\|\gamma\|_{0,p_{0}*}\leq{c_{0}\|\gamma\|_{\varepsilon,p*}},\ \ \gamma\in\mathscr{P}_{\varepsilon,p*}.$
By the upper bound condition in (2.10), we have
$k_{0}:=\frac{p_{0}q}{q-1}\geq k.$
By the definitions of $\eta,p_{0}$ and $k_{0}$, we obtain
$\xi(q):=\delta+\frac{dpq-{k(d+\varepsilon p)(q-1)}}{pqk}=\delta+\frac{d(k_{0}-k)}{k_{0}k}=\eta+\frac{1}{q}\Big(\varepsilon+\frac{d}{p}\Big).$
Then (2.10) implies (2.12), so that
$\eta\leq\xi(q)<1+(2\kappa-\eta)^{+}\leq 1\lor\big(2-i+2\kappa-\eta\big),\ \ i=0,1.$
Moreover, by the second inequality in (2.9), we conclude $\delta+\frac{d}{k}<1\vee(2+2\kappa-\eta)$. Thus, by Lemma 3.4(1), there exists increasing $\beta:(0,\infty)\rightarrow(0,\infty)$ such that
[…]
Next, consider the maximal functional
$\mathscr{M}f(x):=\sup_{r\in(0,1)}\frac{1}{|B(x,r)|}\int_{B(x,r)}f(y)\text{\rm{d}}y,\ \ x\in\mathbb{R}^{d},$
for a nonnegative measurable function $f$. By [18, Lemma 2.1] and $P_{t}^{\gamma}f\in C(\mathbb{R}^{d})$ for $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$ due to (5.5), we find a constant $k_{1}\in(0,\infty)$ such that
$|P_{t}^{\tilde{\gamma}}f(x)-P_{t}^{\tilde{\gamma}}f(y)|\leq k_{1}|x-y|\big(\mathscr{M}|\nabla P_{t}^{\tilde{\gamma}}f|(x)+\mathscr{M}|\nabla P_{t}^{\tilde{\gamma}}f|(y)+{\|P_{t}^{\tilde{\gamma}}f\|_{\infty}}\big),$
$\big\|\mathscr{M}|\nabla P_{t}^{\tilde{\gamma}}f|\big\|_{\tilde{L}^{k_{0}}}\leq k_{1}\|\nabla P_{t}^{\tilde{\gamma}}f\|_{\tilde{L}^{k_{0}}},\ \ \ t\in(0,\tau(\gamma)),\ x,y\in\mathbb{R}^{d}.$
Combining this with Hölder’s inequality, (5.3), (5.4), (5.5) and $\|\gamma\|_{\varepsilon,p*}\geq 1$, we find $k_{2},c_{1}(t)\in(0,\infty)$ such that for any $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$ with $\|f\|_{\tilde{W}^{-\delta,k}}\leq 1$ and $)t\in(0,\tau(\gamma)\land\tau(\tilde{\gamma}$)),
$\big|\gamma(P_{t}^{\tilde{\gamma}}f)-\tilde{\gamma}(P_{t}^{\tilde{\gamma}}f)\big|=\bigg|\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\big(P_{t}^{\tilde{\gamma}}f(x)-P_{t}^{\tilde{\gamma}}f(y)\big)\pi(\text{\rm{d}}x,\text{\rm{d}}y)\bigg|$
$\leq k_{1}\bigg|\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|\big(\mathscr{M}|\nabla P_{t}^{\tilde{\gamma}}f|(x)+\mathscr{M}|\nabla P_{t}^{\tilde{\gamma}}f|(y)+{\|P_{t}^{\tilde{\gamma}}f\|_{\infty}}\big)\pi(\text{\rm{d}}x,\text{\rm{d}}y)\bigg|$
$\leq k_{1}{2^{\frac{1}{q}}}\mathbb{W}_{q}(\gamma,\tilde{\gamma})\Big[(\gamma+\tilde{\gamma})\big((\mathscr{M}|\nabla P_{t}^{\tilde{\gamma}}f|)^{\frac{q}{q-1}}\big)\Big]^{\frac{q-1}{q}}+{k_{1}\mathbb{W}_{1}(\gamma,\tilde{\gamma})\|P_{t}^{\tilde{\gamma}}\|_{\tilde{W}^{-\delta,k}\rightarrow\tilde{W}^{0,\infty}}}$
$\leq k_{1}{2^{\frac{1}{q}}}\mathbb{W}_{q}(\gamma,\tilde{\gamma})(\|\gamma\|_{{p_{0}}*}+\|\tilde{\gamma}\|_{{p_{0}}*})^{\frac{q-1}{q}}\big\|\mathscr{M}|\nabla P_{t}^{\tilde{\gamma}}f|\big\|_{\tilde{L}^{k_{0}}}+{k_{1}\mathbb{W}_{1}(\gamma,\tilde{\gamma})\|P_{t}^{\tilde{\gamma}}\|_{\tilde{W}^{-\delta,k}\rightarrow\tilde{W}^{0,\infty}}}$
$\leq k_{2}\mathbb{W}_{q}(\gamma,\tilde{\gamma})(\|\gamma\|_{\varepsilon,p*}+\|\tilde{\gamma}\|_{\varepsilon,p*})^{\frac{q-1}{q}}\big\|\nabla P_{t}^{\tilde{\gamma}}\big\|_{\tilde{W}^{-\delta,k}\rightarrow\tilde{W}^{0,k_{0}}}+{k_{1}\mathbb{W}_{1}(\gamma,\tilde{\gamma})\|P_{t}^{\tilde{\gamma}}\|_{\tilde{W}^{-\delta,k}\rightarrow\tilde{W}^{0,\infty}}}$
$\leq c_{1}(t)\mathbb{W}_{q}(\gamma,\tilde{\gamma})(\|\gamma\|_{\varepsilon,p*}+\|\tilde{\gamma}\|_{\varepsilon,p*})^{\frac{q-1}{q}}\exp\big[t\beta_{t}k_{t}(\tilde{\gamma})^{\theta}\big]t^{-[(\frac{1+\xi(q)}{2})\vee(\frac{\delta}{2}+\frac{d}{2k})]}.$
Therefore,
[…]
(b) By Duhamel’s formula [8, Proposition 5.5 (2)], we have
$P_{t}^{\gamma}f-P_{t}^{\tilde{\gamma}}f=\int_{0}^{t}P_{s}^{\gamma}\big\langle b_{s}(\cdot,P_{s}^{*}\gamma)-b_{s}(\cdot,P_{s}^{*}\tilde{\gamma}),\ \nabla P_{s,t}^{\tilde{\gamma}}f\big\rangle\text{\rm{d}}s,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d}),$
and (A) implies
$).|b_{t}(x,P_{t}^{*}\gamma)-b_{t}(x,P_{t}^{*}\tilde{\gamma})|\leq K_{t}t^{\kappa}\|P_{t}^{*}\gamma-P_{t}^{*}\tilde{\gamma}\|_{\delta,k*},\ \ t\in[0,\tau(\gamma)\land\tau(\tilde{\gamma}$)).
Then
[…]
where $(\delta^{\prime},k^{\prime})$ is in (4.4) for $r\in[0,1]$ to be determined.
By (2.9), we have either $\eta<\frac{1}{2}+\kappa$ and $\kappa\geq\frac{1}{2}$, or $\eta\in[0,1)$ and $\kappa<\frac{1}{2}$. Let
$r=\begin{cases}1-\kappa,&\text{if}\ \kappa<\frac{1}{2},\eta\in[0,1),\\ \frac{2}{3+2\kappa},&\text{if}\ \eta<\frac{1}{2}+\kappa,\kappa\geq\frac{1}{2}.\end{cases}$
In each case we have
(5.8) $r\eta<1,\ \ \ (2-r)\eta<2+2\kappa,\ \ \ (1-r)\eta\leq 2\kappa,$
where $(2-r)\eta<2+2\kappa$ implies
$\delta^{\prime}-\varepsilon+\frac{d(p-k^{\prime})}{pk^{\prime}}=(1-r)\eta<1\lor(2+2\kappa-\eta),$
so that Lemma 3.4(1) for $i=0$ gives
$\|P_{t}^{\gamma}\|_{\tilde{W}^{-\delta^{\prime},k^{\prime}}\rightarrow\tilde{W}^{-\varepsilon,p}}\leq\beta_{t}\text{\rm{e}}^{t\beta_{t}k_{t}(\gamma)^{\theta}}t^{-\frac{1-r}{2}\eta},$
and by Lemma 3.4(1) for $i=1$,
$\delta-\delta^{\prime}+\frac{d(k^{\prime}-k)}{k^{\prime}k}=r\eta<1$
implies
$\|\nabla P_{s,t}^{\tilde{\gamma}}\|_{\tilde{W}^{-\delta,k}\rightarrow\tilde{W}^{-\delta^{\prime},k^{\prime}}}\leq\beta_{t}\text{\rm{e}}^{t\beta_{t}k_{t}(\tilde{\gamma})^{\theta}}(t-s)^{-\frac{1}{2}-\frac{r}{2}\eta}.$
Combining this with (5.7), we find $c_{2}(t)\in(0,\infty)$ such that
${\big\|P_{t}^{\gamma*}\gamma-P_{t}^{\tilde{\gamma}*}\gamma\big\|_{\delta,k*}}\leq\tilde{D}(t,\gamma,\tilde{\gamma})\int_{0}^{t}s^{\kappa-\frac{1-r}{2}\eta}\big\|P_{s}^{*}\gamma-P_{s}^{*}\tilde{\gamma}\big\|_{\delta,k*}(t-s)^{-\frac{1}{2}-\frac{r}{2}\eta}\text{\rm{d}}s,$
$\tilde{D}(t,\gamma,\tilde{\gamma}):={c_{2}(t)\|\gamma\|_{\varepsilon,p*}\text{\rm{e}}^{t\beta_{t}[k_{t}(\gamma)+k_{t}(\tilde{\gamma})]^{\theta}},\ \ t\in[0,\tau(\gamma)\land\tau(\tilde{\gamma}))}$.
This together with (5.2) and (5.6) yields
$\big\|P_{t}^{*}\gamma-P_{t}^{*}\tilde{\gamma}\big\|_{\delta,k*}\leq{\big\|P_{t}^{*}\gamma-P_{t}^{\tilde{\gamma}*}\gamma\big\|_{\delta,k*}+\big\|P_{t}^{\tilde{\gamma}*}\gamma-P_{t}^{*}\tilde{\gamma}\big\|_{\delta,k*}}$
$\leq D(t,\gamma,\tilde{\gamma})t^{-[(\frac{1+\xi(q)}{2})\vee(\frac{\delta}{2}+\frac{d}{2k})]}+\tilde{D}(t,\gamma,\tilde{\gamma})\int_{0}^{t}s^{\kappa-\frac{1-r}{2}\eta}\big\|P_{s}^{*}\gamma-P_{s}^{*}\tilde{\gamma}\big\|_{\delta,k*}(t-s)^{-\frac{1}{2}-\frac{r}{2}\eta}\text{\rm{d}}s,$
$).\displaystyle\qquad\gamma,\tilde{\gamma}\in\mathscr{P}_{\varepsilon,p*},\ t\in[0,\tau(\gamma)\land\tau(\tilde{\gamma}$)).
[…]
$I_{t}:=\sup_{s\in(0,t]}\text{\rm{e}}^{-\lambda s}s^{[(\frac{1+\xi(q)}{2})\vee(\frac{\delta}{2}+\frac{d}{2k})]}\big\|P_{s}^{*}\gamma-P_{s}^{*}\tilde{\gamma}\big\|_{\delta,k*}<\infty$
and
$I_{t}\leq$ $\,D(t,\gamma,\tilde{\gamma})+\tilde{D}(t,\gamma,\tilde{\gamma})\sup_{s\in(0,t]}s^{[(\frac{1+\xi(q)}{2})\vee(\frac{\delta}{2}+\frac{d}{2k})]+([(\frac{1+\xi(q)}{2})\vee(\frac{\delta}{2}+\frac{d}{2k})]+\frac{1-r}{2}\eta-\kappa)^{-}}$
$\qquad\qquad\qquad\quad\quad\times\int_{0}^{s}u^{-([(\frac{1+\xi(q)}{2})\vee(\frac{\delta}{2}+\frac{d}{2k})]+\frac{1-r}{2}\eta-\kappa)^{+}}\text{\rm{e}}^{-\lambda(s-u)}(s-u)^{-\frac{1}{2}-\frac{r}{2}\eta}\text{\rm{d}}u.$
By (5.8), (2.12) and $\delta+\frac{d}{k}<2+(2\kappa-\eta)^{+}$, we have
$(\frac{1+\xi(q)}{2})\vee(\frac{\delta}{2}+\frac{d}{2k})+\frac{1-r}{2}\eta-\kappa<1,\ \ \ \frac{1}{2}+\frac{r}{2}\eta<1,$
we may apply (3.1) to $\alpha_{1}=([(\frac{1+\xi(q)}{2})\vee(\frac{\delta}{2}+\frac{d}{2k})]+\frac{1-r}{2}\eta-\kappa)^{+},\alpha_{2}=\frac{1}{2}+\frac{r}{2}\eta$ and $\alpha=\frac{1}{\theta},$ such that
$I_{t}\leq D(t,\gamma,\tilde{\gamma})+c\tilde{D}(t,\gamma,\tilde{\gamma})\lambda^{-\frac{1}{\theta}}$
holds for some constant $c\in(0,\infty)$. Taking
$\lambda=\big(2c\tilde{D}(t,\gamma,\tilde{\gamma})\big)^{\theta}$
we obtain
[…]
which implies (2.11) for some increasing $\beta:(0,\infty)\rightarrow(0,\infty).$
Finally, when $\varepsilon=0,p=\infty$, $k_{t}(\gamma)$ defined in (2.7) is bounded above by some constant $c(t)\in(0,\infty)$ uniformly in $\gamma\in\mathscr{P}_{\varepsilon,p*}=\mathscr{P}$. Moreover, in this case (2.9) implies that $q=1$ satisfies (2.10). So, (2.11) implies (2.13). ∎
Proof of Theorem 2.3(2).
The proof is similar to that of [8, Theorem 2.3(2)], where $\varepsilon=\delta=0.$ For completeness we figure out it in the present situation.
(a) For $)t\in(0,\tau(\gamma)\land\tau(\tilde{\gamma}$)), we consider the SDEs
(5.9) $\begin{split}&\text{\rm{d}}X_{s}=b_{s}(X_{s},P_{s}^{*}\gamma)\text{\rm{d}}s+\text{\rm{d}}W_{s},\\ &\text{\rm{d}}Y_{s}=b_{s}(Y_{s},P_{s}^{*}\tilde{\gamma})\text{\rm{d}}s+\text{\rm{d}}W_{s},\ \ \ s\in[0,t],\end{split}$
where $X_{0},Y_{0}$ are $\mathscr{F}_{0}$-measurable such that
$\mathscr{L}_{X_{0}}=\gamma,\ \ \ \mathscr{L}_{Y_{0}}=\tilde{\gamma},\ \ \ \mathbb{E}|X_{0}-Y_{0}|^{2}=\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2}.$
Then
(5.10) $P_{t}^{*}\gamma=\mathscr{L}_{X_{t}},\ \ \ \ P_{t}^{*}\tilde{\gamma}=\mathscr{L}_{Y_{t}}.$
[…]
For fixed $\theta^{\prime}\in(\theta,\infty)$, let
[…]
According to the bi-coupling method developed in [12], the following SDE will be coupled with those two SDEs in (5.9) respectively:
$\text{\rm{d}}Z_{s}=\Big(1_{[0,t_{0}]}(s)b_{s}(Z_{s},P_{s}^{\ast}\tilde{\gamma})+1_{(t_{0},t]}(s)b_{s}(Z_{s},P_{s}^{\ast}\gamma)\Big)\text{\rm{d}}s+\text{\rm{d}}W_{s},\ \ Z_{0}=Y_{0},s\in[0,t].$
By (5.10) and [12, Lemma 2.1], we have
(5.12) $\begin{split}&{\rm Ent}(P_{t}^{\ast}\gamma|P_{t}^{\ast}\tilde{\gamma})={\rm Ent}(\mathscr{L}_{X_{t}}|\mathscr{L}_{Y_{t}})\\ &\leq 2{\rm Ent}(\mathscr{L}_{X_{t}}|\mathscr{L}_{Z_{t}})+\log\int_{\mathbb{R}^{d}}\Big(\frac{\text{\rm{d}}\mathscr{L}_{Z_{t}}}{\text{\rm{d}}\mathscr{L}_{Y_{t}}}\Big)^{2}\text{\rm{d}}\mathscr{L}_{Y_{t}}=:2I_{1}+I_{2}.\end{split}$
Below we estimate $I_{1}$ and $I_{2}$ respectively.
(b) To estimate $I_{1}$, we first establish the log-Harnack inequality for $P_{t}^{\gamma}:$ for any $\theta^{\prime}\in(\theta,\infty)$, there exists $c_{0}(t)\in(0,\infty)$ increasingly in $t$ such that
(5.13) $\begin{split}&P_{s,t}^{\gamma}\log f(x)\leq\log P_{s,t}^{\gamma}f(y)+\frac{c_{0}(t)|x-y|^{2}}{s_{t}(\theta^{\prime},\gamma)\wedge(t-s)},\\ &\qquad\ x,y\in\mathbb{R}^{d},0\leq s<t<\tau(\gamma),\ \gamma\in\mathscr{P}_{\varepsilon,p*},\ f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{d})\end{split}$
for $s_{t}(\theta^{\prime},\gamma)$ defined in (2.8). By (A), we have
$\|b_{t}(\cdot,P_{t}^{\ast}\gamma)\|_{\infty}\leq K_{t}k_{t}(\gamma)t^{\kappa-\frac{\eta}{2}},\ \ \ t\in[0,\tau(\gamma)).$
Since (2.9) implies (2.1) so that $(\eta-2\kappa)^{+}<1$. For any constant
$\theta^{\prime}>\theta:=\frac{2}{1-(\eta-2\kappa)^{+}},$
we have
$q^{\prime}:={\Big(\frac{(\eta-2\kappa)^{+}}{2}+\frac{1}{\theta^{\prime}}\Big)^{-1}}\in\Big(2,\frac{2}{(\eta-2\kappa)^{+}}\Big),$
and for some $c_{1}(t)\in(0,\infty)$ increasing in $t$,
(5.14) $\|b_{\cdot}(\cdot,\gamma_{\cdot})\|_{\tilde{L}_{q^{\prime}}^{\infty}(s,t)}\leq c_{1}(t)k_{t}(\gamma)(t-s)^{\frac{1}{\theta^{\prime}}},\ \ 0\leq s<t<\tau(\gamma).$
By (2.8), we find a constant $k_{1}\in(0,\infty)$ such that
$k_{t}(\gamma)(t-s)^{\frac{1}{\theta^{\prime}}}\leq k_{1},\ \ 0<t-s\leq s_{t}(\theta^{\prime},\gamma).$
This together with (5.14) implies that
$\|b_{\cdot}(\cdot,\gamma_{\cdot})\|_{\tilde{L}_{q^{\prime}}^{\infty}(s,t)}\leq{k_{1}c_{1}(t)},\ \ 0<t-s\leq s_{t}(\theta^{\prime},\gamma),\ t\in(0,\tau(\gamma)).$
So, by [8, Proposition 5.2(4)], we derive (5.13) for $0<t-s\leq s_{t}(\theta^{\prime},\gamma),\ t\in(0,\tau(\gamma))$ and some $c_{0}(t)\in(0,\infty)$ increasingly in $t$. When $t-s>s_{t}(\theta^{\prime},\gamma),\ t\in(0,\tau(\gamma)),$ by the semigroup property and Jensen’s inequality, we deduce
$P_{s,t}^{\gamma}\log f(x)=P_{s,s+s_{t}(\theta^{\prime},\gamma)}^{\gamma}P_{s+s_{t}(\theta^{\prime},\gamma),t}^{\gamma}\log f(x)\leq P_{s,s+s_{t}(\theta^{\prime},\gamma)}^{\gamma}\log P_{s+s_{t}(\theta^{\prime},\gamma),t}^{\gamma}f(x)$
$\leq\log P_{s,s+s_{t}(\theta^{\prime},\gamma)}^{\gamma}P_{s+s_{t}(\theta^{\prime},\gamma),t}^{\gamma}f(y)+\frac{{c_{0}(t)|x-y|^{2}}}{s_{t}(\theta^{\prime},\gamma)}$
$=\log P_{s,t}^{\gamma}f(y)+\frac{c_{0}(t)|x-y|^{2}}{s_{t}(\theta^{\prime},\gamma)},\ \ x,y\in\mathbb{R}^{d},\ f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{d}).$
So, (5.13) also holds for $t-s>s_{t}(\theta^{\prime},\gamma)$.
Next, by the Markov property, we have
$\mathbb{E}[f(X_{t})]=\mathbb{E}[(P_{t_{0},t}^{\gamma}f)(X_{t_{0}})],\ \ \ \mathbb{E}[f(Z_{t})]=\mathbb{E}[(P_{t_{0},t}^{\gamma}f)(Z_{t_{0}})].$
This together with (5.13) for $s=t_{0}$ and Jensen’s inequality implies
(5.15) $\mathbb{E}[\log f(X_{t})]\leq\log\mathbb{E}[f(Z_{t})]+\frac{2c_{0}(t)}{s_{t}(\theta^{\prime},\gamma)}\mathbb{E}[|X_{t_{0}}-Z_{t_{0}}|^{2}],\ f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{d}).$
Moreover, by (A), (2.11) and (5.11), we find $c_{2}(t)\in(0,\infty)$ increasing in $t$ such that
$\int_{0}^{t_{0}}\|b_{s}(\cdot,P_{s}^{\ast}\gamma)-b_{s}(\cdot,P_{s}^{\ast}\tilde{\gamma})\|_{\infty}\text{\rm{d}}s\leq K_{t}\int_{0}^{t_{0}}\|P_{s}^{\ast}\gamma-P_{s}^{\ast}\tilde{\gamma}\|_{\delta,k}s^{\kappa}\text{\rm{d}}s$
$\leq K_{t}\big(\|\gamma\|_{\varepsilon,p*}+\|\tilde{\gamma}\|_{\varepsilon,p*}\big)^{\frac{q-1}{q}}K_{t,\beta}^{(\theta)}(\gamma,\tilde{\gamma})\mathbb{W}_{q}(\gamma,\tilde{\gamma})\int_{0}^{t_{0}}s^{\kappa-[(\frac{1+\xi(q)}{2})\vee(\frac{\delta}{2}+\frac{d}{2k})]}\text{\rm{d}}s$
$\leq c_{2}(t)\big(\|\gamma\|_{\varepsilon,p*}+\|\tilde{\gamma}\|_{\varepsilon,p*}\big)^{\frac{q-1}{q}}\mathbb{W}_{q}(\gamma,\tilde{\gamma})t_{0}^{{(1+\kappa)-[(\frac{1+\xi(q)}{2})\vee(\frac{\delta}{2}+\frac{d}{2k})]}}.$
Combining this with [8, Proposition 5.5(1)], we find $c_{2}(t)\in(0,\infty)$ increasing in $t$ such that
$\mathbb{E}[|X_{t_{0}}-Z_{t_{0}}|^{2}]\leq c_{2}(t)\big(\|\gamma\|_{\varepsilon,p*}+\|\tilde{\gamma}\|_{\varepsilon,p*}\big)^{\frac{2(q-1)}{q}}\big(\mathbb{W}_{q}(\gamma,\tilde{\gamma})^{2}{t_{0}}^{2(1+\kappa)-[(1+\xi(q))\vee(\delta+\frac{d}{k})]}+\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2}\big).$
This together with the formula
${\rm Ent}(\mu|\nu)=\sup_{f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{d})}\big(\mu(\log f)-\log\nu(f)\big),\ \ \mu,\nu\in\mathscr{P}$
and (5.15) yields that for some constant $c_{3}(t)\in(0,\infty)$ increasing in $t$ such that
(5.16) $\begin{split}&I_{1}:={\rm Ent}(\mathscr{L}_{X_{t}}|\mathscr{L}_{Z_{t}})=\sup_{f\in\mathscr{B}_{b}^{+}(\mathbb{R}^{d})}\big(\mathbb{E}[\log f(X_{t})]-\log\mathbb{E}[f(Z_{t})]\big),\\ &\leq c_{3}(t)(\|\gamma\|_{\varepsilon,p*}\lor\|\tilde{\gamma}\|_{\varepsilon,p*})^{\frac{2(q-1)}{q}}\bigg(\frac{\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2}}{s_{t}(\theta^{\prime},\gamma)}+\frac{\mathbb{W}_{q}(\gamma,\tilde{\gamma})^{2}}{s_{t}(\theta^{\prime},\gamma)^{([(1+\xi(q))\vee(\delta+\frac{d}{k})]-(2\kappa+1))^{+}}}\bigg),\end{split}$
where in the last step we have used $t_{0}\leq s_{t}(\theta^{\prime},\gamma)$.
(c) Estimate $I_{2}$. By (A) and (2.11), we find a constant $K(t)\in(0,\infty)$ increasing in $t$ such that
[…]
satisfies
(5.17) $\begin{split}&|\xi_{s}|^{2}\leq K_{t}^{2}C(t,\gamma,\tilde{\gamma})s^{2\kappa-[(1+\xi(q))\vee(\delta+\frac{d}{k})]},\ \ \ s\in(0,t],\\ &C(t,\gamma,\tilde{\gamma}):=(\|\gamma\|_{\varepsilon,p*}+\|\tilde{\gamma}\|_{\varepsilon,p*})^{\frac{2(q-1)}{q}}K_{t,\beta}^{(\theta)}(\gamma,\tilde{\gamma})^{2}\mathbb{W}_{q}(\gamma,\tilde{\gamma})^{2}.\end{split}$
Then
$R_{s}:=\text{\rm{e}}^{\int_{t_{0}}^{s}\langle\xi_{r},\text{\rm{d}}W_{r}\rangle-\frac{1}{2}\int_{t_{0}}^{s}|\xi_{r}|^{2}\text{\rm{d}}r},\ \ s\in[t_{0},t]$
is a martingale, and by Girsanov’s theorem,
$\frac{\text{\rm{d}}\mathscr{L}_{Z_{t}}}{\text{\rm{d}}\mathscr{L}_{Y_{t}}}(Y_{t})=\mathbb{E}(R_{t}|Y_{t}).$
Combining this with Jensen’s inequality and (5.17), we find $c_{4}(t)\in(0,\infty)$ increasing in $t$ such that
$I_{2}$ $:=\log\mathbb{E}\bigg[\Big(\frac{\text{\rm{d}}\mathscr{L}_{Z_{t}}}{\text{\rm{d}}\mathscr{L}_{Y_{t}}}(Y_{t})\Big)^{2}\bigg]\leq\log\mathbb{E}\Big[R_{t}^{2}\Big]$
$\leq\log\mathbb{E}\bigg[\text{\rm{e}}^{2\int_{t_{0}}^{t}\langle\xi_{s},\text{\rm{d}}W_{s}\rangle-2\int_{t_{0}}^{t}|\xi_{s}|^{2}\text{\rm{d}}s+C(t,\gamma,\tilde{\gamma})\int_{t_{0}}^{t}s^{2\kappa-1-\xi(q)}\text{\rm{d}}s}\bigg]$
$=C(t,\gamma,\tilde{\gamma})\int_{t_{0}}^{t}s^{2\kappa-[(1+\xi(q))\vee(\delta+\frac{d}{k})]}\text{\rm{d}}s\leq c_{3}(t)C(t,\gamma,\tilde{\gamma})t_{0}^{-([(1+\xi(q))\vee(\delta+\frac{d}{k})]-(2\kappa+1))^{+}}.$
By combining this with (5.12) and (5.16), we obtain (2.14) for some $\beta:(0,\infty)\rightarrow(0,\infty).$
(c) If $\varepsilon=0,p=\infty$, we have $\mathscr{P}_{\varepsilon,p*}=\mathscr{P}$, $\tau(\gamma)=\infty$ and $\|\gamma\|_{\varepsilon,p*}=1$ for any $\gamma\in\mathscr{P}$, and we may take $q=1$ so that $\xi(q)=\eta=\delta+\frac{d}{k}$, and (2.12) implies
$(1+\xi(q))\vee(\delta+\frac{d}{k})]-(2\kappa+1)<1.$
Hence (2.14) implies (2.15) for some increasing $\beta:(0,\infty)\rightarrow(0,\infty).$

### v2

@@
 Moreover, the relative entropy and the $\|\cdot\|_{\delta,k*}$-distance induced by $\tilde{W}^{-\delta,k}$ are estimated for the time-marginal distributions of solutions by using the Wasserstein distance of initial distributions, which describe the regularity of the solution in initial distribution.
-In particular, the main results apply to Nemytskii-type SDEs which point-wisely depend on the density function and its derivatives, as well as McKean-Vlasov SDEs with interactions more singular than Riesz kernels.
+In particular, the main results apply to Nemytskii-type SDEs depending on higher order derivatives of density functions, as well as McKean-Vlasov SDEs with interactions more singular than Riesz kernels.
 AMS Subject Classification: 60H10, 60H50.
-Keywords: McKean-Vlasov SDE, local distributional interaction, $\|\cdot\|_{\delta,k*}$-distance,
-entropy-cost inequality.
+Keywords: McKean-Vlasov SDE, local distributional interaction, well-posedness,
+$\|\cdot\|_{\delta,k*}$-distance, entropy-cost inequality.
 1 Introduction
@@
 for $B(z,1):=\{x\in\mathbb{R}^{d}:|x-z|\leq 1\}$ and $\|\cdot\|_{L^{k}}$ being the $L^{k}$-norm with respect to the Lebesgue measure.
-Let $\tilde{W}_{*}^{-\delta,k}$ be the dual space of $\tilde{W}^{-\delta,k}.$ We study (1.1) for $h_{t}\in\tilde{W}^{-\delta,k}$ for some $(\delta,k)\in[0,\infty)\times[1,\infty],$ where $b_{t}$ in (1.2) is well-defined as duality for
-$\mu\in\mathscr{P}_{\delta,k*}:=\mathscr{P}\cap\tilde{W}_{*}^{-\delta,k}:=\left\{\mu\in\mathscr{P}:\ \|\mu\|_{\delta,k*}:=\sup_{f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\|f\|_{\tilde{W}^{-\delta,k}}\leq 1}|\mu(f)|<\infty\right\}.$
+Let $\tilde{W}_{*}^{-\delta,k}$ be the dual space of $\tilde{W}^{-\delta,k}.$ We study (1.1) for $h_{t}\in\tilde{W}^{-\delta,k}$ for some $(\delta,k)\in[0,\infty)\times[1,\infty],$ where $b_{t}$ in (1.2) is well-defined as duality for $\mu\in\mathscr{P}_{\delta,k*}:=\mathscr{P}\cap\tilde{W}_{*}^{-\delta,k}.$ We have
+$\mathscr{P}_{\delta,k*}=\left\{\mu\in\mathscr{P}:\ \|\mu\|_{\delta,k*}:=\sup_{f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\|f\|_{\tilde{W}^{-\delta,k}}\leq 1}|\mu(f)|<\infty\right\}.$
 It is easy to see that $\mathscr{P}_{\delta,k*}$ is a closed subset of $\tilde{W}^{-\delta,k}_{*},$ where $\tilde{W}^{-\delta,k}_{*}$ is a Banach space with norm
@@
 2 Main results
+2.1 Well-posedness
 To solve (1.1), we make the following assumptions where the drift is Lipschitz continuous in distribution under the $\|\cdot\|_{\delta,k*}$ distance.
@@
 Then assertions in Theorem 2.1 hold for $\varepsilon\in[0,\delta]$ and $p\in[k,\infty]$ satisfying (2.1).
-When $\kappa=0$, the condition (2.1) coincides with [4, (1.33)] for $\alpha=2,q_{b}=\infty,p_{0}=\frac{p}{p-1},\rho_{0}=k,\beta_{0}=\varepsilon$ and $\beta_{b}=-\delta$.
-In this case, [4, Theorem 1.9] ensures the well-posedness and density estimates for (1.1) provided $\|h_{t}\|_{\mathbb{B}_{k,\infty}^{\delta}}<\infty$.
-If moreover $p=\infty$ and $\varepsilon=0$, (2.1) becomes $\delta<1-\frac{d}{k},$ so that [1, Theorem 1] with $\alpha=2$ implies the global well-posedness of (1.1) for $\|h\|_{\mathbb{B}_{k,\infty}^{-\delta}}<\infty.$ These results do not cover Theorem 2.1 since $\|h\|_{\mathbb{B}_{k,\infty}^{-\delta}}<\infty$ may fail for $h\in\tilde{W}^{-\delta,k}(\mathbb{R}^{d};\mathbb{R}^{d}).$
+When $\kappa=0$, the condition (2.1) coincides with [6, (1.33)] for $\alpha=2,q_{b}=\infty,p_{0}=\frac{p}{p-1},\rho_{0}=k,\beta_{0}=\varepsilon$ and $\beta_{b}=-\delta$.
+In this case, [6, Theorem 1.9] ensures the well-posedness and density estimates for (1.1) provided $\|h_{t}\|_{\mathbb{B}_{k,\infty}^{\delta}}<\infty$.
+If moreover $p=\infty$ and $\varepsilon=0$, (2.1) becomes $\delta<1-\frac{d}{k},$ so that [2, Theorem 1] with $\alpha=2$ implies the global well-posedness of (1.1) for $\|h\|_{\mathbb{B}_{k,\infty}^{-\delta}}<\infty.$ These results do not cover Theorem 2.1 since $\|h\|_{\mathbb{B}_{k,\infty}^{-\delta}}<\infty$ may fail for $h\in\tilde{W}^{-\delta,k}(\mathbb{R}^{d};\mathbb{R}^{d}).$
 For $\delta\in[0,\infty)$ and $k\in[1,\infty]$, let $W^{-\delta,k}(\mathbb{R}^{d};\mathbb{R}^{d})$ be the closure of $C_{0}^{\infty}(\mathbb{R}^{d};\mathbb{R}^{d})$ with respect to the negative Sobolev norm $\|f\|_{W^{-\delta,k}}:=\|(1-\Delta)^{-\frac{\delta}{2}}f\|_{L^{k}}$.
@@
 There are also many other papers concerning log-Harnack inequalities and applications, which we do not mention in details to save space.
-We present the following two examples to illustrate Theorem 2.1 and Theorem 2.3, where the kernel $h$ is more singular than the Riesz kernel as considered in previous papers, see [8, 11, 13, 14] and references therein, and covers the Dirac interaction for Nemytskii-type SDEs depending on the density and any order derivatives.
+We present the following two examples to illustrate Theorem 2.1 and Theorem 2.3, where the kernel $h$ is more singular than the Riesz kernel as considered in previous papers, see [5, 10, 13, 15, 16] and references therein.
+In particular, our results apply to the Dirac interaction for Nemytskii-type SDEs depending on the density and its higher order derivatives, which include the Burgers/Navier-Stokes/$p$-Laplacian equations as typical examples where the first and second order derivatives of density are involved in, see for instance [1, 17].
 Example 2.5 (Interactions with Arbitrary Singular Order).
@@
 $|(1-\Delta)^{n}\nabla^{i}P_{t}^{0}f|\leq C_{n+i}t^{-\frac{i}{2}-n}P_{2t}^{0}|f|,\ \ t>0,\ i,n\in\mathbb{Z}_{+},f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
-On the other hand, by [8, Lemma 5.3], there exists a constant $c_{0}\in(0,\infty)$ such that
+On the other hand, by [10, Lemma 5.3], there exists a constant $c_{0}\in(0,\infty)$ such that
 $\|P_{2t}^{0}\|_{\tilde{L}^{k}\rightarrow\tilde{L}^{p}}\leq c_{0}t^{-\frac{d(p-k)}{2pk}},\ \ t>0,\ 1\leq k\leq p\leq\infty.$
@@
 (3.6) $\|b^{\mu}\|_{\tilde{L}_{q^{\prime}}^{\infty}(T)}=\sup_{z\in\mathbb{R}^{d}}\left(\int_{0}^{T}\|b_{t}^{\mu}1_{B(z,1)}\|_{\infty}^{q^{\prime}}\right)^{\frac{1}{q^{\prime}}}<\infty.$
-Then the desired assertion follows from [8, Proposition 5.1]. ∎
+Then the desired assertion follows from [10, Proposition 5.1]. ∎
 For $\mu\in\mathscr{C}_{\varepsilon,p;\delta,k}^{T}$, under (2.1) we denote
@@
 (3.10) $\sup_{0\leq r\leq s\leq t}(s-r)^{\frac{i}{2}}\|\nabla^{i}{P}_{r,s}^{\mu}\|_{\tilde{W}^{-\varepsilon_{2},p_{2}}\rightarrow\tilde{W}^{-\varepsilon_{2},p_{2}}}<\infty,\ \ i=0,1.$
-By [8, Proposition 5.4], (3.6) implies that for $p_{2}>1$
+By [10, Proposition 5.4], (3.6) implies that for $p_{2}>1$
 (3.11) $\sup_{0\leq r\leq s\leq t}(s-r)^{\frac{i}{2}}\|\nabla^{i}P_{r,s}^{\mu}\|_{\tilde{L}^{p_{2}}\rightarrow\tilde{L}^{p_{2}}}<\infty,\ \ i=0,1.$
-Moreover, by the Duhamel formula, see [8, Proposition 5.5(2)], we have
+Moreover, by the Duhamel formula, see [10, Proposition 5.5(2)], we have
 (3.12) ${P}_{s,t}^{\mu}f=P_{t-s}^{0}f+\int_{s}^{t}{P}_{s,r}^{\mu}\langle b_{r}(\cdot,\mu_{r}),\nabla P_{t-r}^{0}f\rangle\text{\rm{d}}r,\ \ s\in[0,t],\ f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
@@
 by (3.1) with $\alpha=0$, we find $\beta_{l+1}(t)\in(0,\infty)$ increasing in $t>0$ such that (3.16) holds for $l+1$ in place $l$. Then the proof is finished. ∎
-4 Proof of Theorem 2.1
+4 Proof of the existence and uniqueness
+In this section we prove Theorem 2.1.
 Let $X_{0}$ be $\mathscr{F}_{0}$-measurable such that $\gamma:=\mathscr{L}_{X_{0}}\in\mathscr{P}_{\varepsilon,p*},$ and let $T\in(0,\infty)$ be fixed.
@@
 Proof.
-According to the proof of [8, Theorem 2.1], the first assrtion follows from the second.
+According to the proof of [10, Theorem 2.1], the first assrtion follows from the second.
 So, it suffices to show that for any $n\in\mathbb{N}$ and initial value $X_{0}$ with $\gamma\in\mathscr{L}_{X_{0}}\in\mathscr{P}_{\varepsilon,p*},$ (1.1) has a unique weak/strong $\mathscr{C}_{\varepsilon,p;\delta,k}$-solution up to $\tau_{n}(\gamma)$ such that
@@
 Then by (2.1) and (2.5), there exists $q^{\prime}>2$ such that $\|b^{\mu}\|_{\tilde{L}_{q^{\prime}}^{\infty}(\tau_{n})}$ in (3.6) is bounded above by a constant $D_{n,\gamma}\in(0,\infty)$ depending on $n$ and $\|\gamma\|_{\varepsilon,p*}$.
-So, (4.15) follows from [17, Theorem 1.3.1], where $c_{n,q}(\gamma)$ is uniform in $\gamma\in\mathscr{P}$ when $\varepsilon=0$ and $p=\infty$, since $\|\gamma\|_{0,\infty*}=1$.
+So, (4.15) follows from [20, Theorem 1.3.1], where $c_{n,q}(\gamma)$ is uniform in $\gamma\in\mathscr{P}$ when $\varepsilon=0$ and $p=\infty$, since $\|\gamma\|_{0,\infty*}=1$.
 ∎
@@
 ∎
-5 Proof of Theorem 2.3
+5 Proof of regularity estimates
 According to Theorem 2.1(3), Theorem 2.3(3) follows from Theorem 2.3(1)-(2). So, we only prove the first two assertions in Theorem 2.3.
@@
 for a nonnegative measurable function $f$.
-By [18, Lemma 2.1] and $P_{t}^{\gamma}f\in C(\mathbb{R}^{d})$ for $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$ due to (5.5), we find a constant $k_{1}\in(0,\infty)$ such that
+By [21, Lemma 2.1] and $P_{t}^{\gamma}f\in C(\mathbb{R}^{d})$ for $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$ due to (5.5), we find a constant $k_{1}\in(0,\infty)$ such that
 $|P_{t}^{\tilde{\gamma}}f(x)-P_{t}^{\tilde{\gamma}}f(y)|\leq k_{1}|x-y|\big(\mathscr{M}|\nabla P_{t}^{\tilde{\gamma}}f|(x)+\mathscr{M}|\nabla P_{t}^{\tilde{\gamma}}f|(y)+{\|P_{t}^{\tilde{\gamma}}f\|_{\infty}}\big),$
@@
 […]
-(b) By Duhamel’s formula [8, Proposition 5.5 (2)], we have
+(b) By Duhamel’s formula [10, Proposition 5.5 (2)], we have
 $P_{t}^{\gamma}f-P_{t}^{\tilde{\gamma}}f=\int_{0}^{t}P_{s}^{\gamma}\big\langle b_{s}(\cdot,P_{s}^{*}\gamma)-b_{s}(\cdot,P_{s}^{*}\tilde{\gamma}),\ \nabla P_{s,t}^{\tilde{\gamma}}f\big\rangle\text{\rm{d}}s,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d}),$
@@
 Proof of Theorem 2.3(2).
-The proof is similar to that of [8, Theorem 2.3(2)], where $\varepsilon=\delta=0.$ For completeness we figure out it in the present situation.
+The proof is similar to that of [10, Theorem 2.3(2)], where $\varepsilon=\delta=0.$ For completeness we figure out it in the present situation.
 (a) For $)t\in(0,\tau(\gamma)\land\tau(\tilde{\gamma}$)), we consider the SDEs
@@
 $\text{\rm{d}}Z_{s}=\Big(1_{[0,t_{0}]}(s)b_{s}(Z_{s},P_{s}^{\ast}\tilde{\gamma})+1_{(t_{0},t]}(s)b_{s}(Z_{s},P_{s}^{\ast}\gamma)\Big)\text{\rm{d}}s+\text{\rm{d}}W_{s},\ \ Z_{0}=Y_{0},s\in[0,t].$
-By (5.10) and [12, Lemma 2.1], we have
+By (5.10) and [14, Lemma 2.1], we have
 (5.12) $\begin{split}&{\rm Ent}(P_{t}^{\ast}\gamma|P_{t}^{\ast}\tilde{\gamma})={\rm Ent}(\mathscr{L}_{X_{t}}|\mathscr{L}_{Y_{t}})\\ &\leq 2{\rm Ent}(\mathscr{L}_{X_{t}}|\mathscr{L}_{Z_{t}})+\log\int_{\mathbb{R}^{d}}\Big(\frac{\text{\rm{d}}\mathscr{L}_{Z_{t}}}{\text{\rm{d}}\mathscr{L}_{Y_{t}}}\Big)^{2}\text{\rm{d}}\mathscr{L}_{Y_{t}}=:2I_{1}+I_{2}.\end{split}$
@@
 $\|b_{\cdot}(\cdot,\gamma_{\cdot})\|_{\tilde{L}_{q^{\prime}}^{\infty}(s,t)}\leq{k_{1}c_{1}(t)},\ \ 0<t-s\leq s_{t}(\theta^{\prime},\gamma),\ t\in(0,\tau(\gamma)).$
-So, by [8, Proposition 5.2(4)], we derive (5.13) for $0<t-s\leq s_{t}(\theta^{\prime},\gamma),\ t\in(0,\tau(\gamma))$ and some $c_{0}(t)\in(0,\infty)$ increasingly in $t$.
+So, by [10, Proposition 5.2(4)], we derive (5.13) for $0<t-s\leq s_{t}(\theta^{\prime},\gamma),\ t\in(0,\tau(\gamma))$ and some $c_{0}(t)\in(0,\infty)$ increasingly in $t$.
 When $t-s>s_{t}(\theta^{\prime},\gamma),\ t\in(0,\tau(\gamma)),$ by the semigroup property and Jensen’s inequality, we deduce
@@
 $\leq c_{2}(t)\big(\|\gamma\|_{\varepsilon,p*}+\|\tilde{\gamma}\|_{\varepsilon,p*}\big)^{\frac{q-1}{q}}\mathbb{W}_{q}(\gamma,\tilde{\gamma})t_{0}^{{(1+\kappa)-[(\frac{1+\xi(q)}{2})\vee(\frac{\delta}{2}+\frac{d}{2k})]}}.$
-Combining this with [8, Proposition 5.5(1)], we find $c_{2}(t)\in(0,\infty)$ increasing in $t$ such that
+Combining this with [10, Proposition 5.5(1)], we find $c_{2}(t)\in(0,\infty)$ increasing in $t$ such that
 $\mathbb{E}[|X_{t_{0}}-Z_{t_{0}}|^{2}]\leq c_{2}(t)\big(\|\gamma\|_{\varepsilon,p*}+\|\tilde{\gamma}\|_{\varepsilon,p*}\big)^{\frac{2(q-1)}{q}}\big(\mathbb{W}_{q}(\gamma,\tilde{\gamma})^{2}{t_{0}}^{2(1+\kappa)-[(1+\xi(q))\vee(\delta+\frac{d}{k})]}+\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2}\big).$
@@
 Hence (2.14) implies (2.15) for some increasing $\beta:(0,\infty)\rightarrow(0,\infty).$
+Data Availability
+Data sharing is not applicable to this paper as no datasets were generated or analyzed.
+Declarations
+The authors declare that they have no Conflict of interests.